### v1

# A Mind Cannot Be Smeared Across Time

Abstract
Whether machines can be conscious depends not only on what they compute, but when they compute it. Most deployed artificial systems realise their functions via sequential or time-multiplexed updates. Conscious experience appears unified and simultaneous. I show that this difference matters formally. I augment Stack Theory with algebraic laws relating within time-window constraint satisfaction to conjunction. I introduce a precise temporal semantics over windowed trajectories $\tau^{\Delta,s}$ and prove that existential temporal realisation $\Diamond_{\Delta}$ does not preserve conjunction. A system can realise all the ingredients of experience across time without ever instantiating the experienced conjunction itself. I then distinguish two postulates. StrongSync requires objective co-instantiation of the grounded conjunction within the window, while WeakSync permits temporal “smearing”. I formalise concurrency-capacity to measure what is needed to satisfy StrongSync. Finally, I review neurophysiological evidence suggesting that consciousness depends on phase synchrony and effective connectivity, and that loss of consciousness is often associated with its breakdown. This evidence makes WeakSync less plausible. Under StrongSync, software consciousness on strictly sequential substrates is impossible for contents whose grounding requires two or more simultaneous contributors. The more parts from which simultaneous contribution required, the more concurrency capacity is required. The hardware matters. Consciousness attribution therefore requires architectural inspection, not just functional performance.
1 Introduction
A central open question in machine consciousness is whether behaviourally and functionally matched systems must be matched in phenomenal realisation. Stack Theory seeks to approach this question by formalising cognitive processes in terms of abstraction layers [undeff, undefb, undefh]. Various aspects of incumbent theories of consciousness are formalised [undefw, undefj, undefn, undefv, undefy, undef], organised into a hierarchy of causal identities [undefb], and derived as a necessary consequence of a generalisation-optimal learning theory that makes valence primary [undefe, undefg, undefc, undefa]. Those necessary ingredients of consciousness are beyond the scope of this paper. Here I am concerned with a question of sufficiency. Crucially, Stack Theory shows how the behaviour of two systems can appear identical at one level of abstraction, while being different in the levels below. This means that conscious experience can be purely functional, but still depend on more than surface level function. It can depend not just on what is computed, but when and how. The ingredients of what looks like a single experienced moment at a high level of abstraction may be smeared across objective micro-time at a lower level of abstraction, such that no objective time-slice contains the full grounded conjunction (meaning the parts of a subjective moment of consciousness exist at different objective times, never together). If the ingredients of a subjective moment need not synchronise in objective time for that moment to be experienced, then current hardware is sufficient to host machine consciousness. That also means a “liquid-brained” swarm intelligence [undefx] like an ant colony or a population of humans can be conscious. However, if synchronised broadcast and causal co-dependence matter, then far fewer systems can be conscious. The hardware matters, because the parts of a subjective conscious moment need to be connected in objective time to be experienced. This question of architectural sufficiency is called The Temporal Gap.
The Temporal Gap was proposed in earlier work [undeff], but was not rigorously formalised. Here I seek to rectify that oversight with two kinds of contribution. Formally, I add a Stack-Time Semantics module to Stack Theory. This module includes induced layer time, window environments, temporal lifting operators, and algebraic laws relating “within-window” satisfaction to conjunction. Using this, I refine the Temporal Gap into a testable architectural question about synchrony and concurrency.
Relation to existing consciousness theories.
Beyond Stack Theory, there are many scientific theories which emphasise some form of integration or unity across time and causal organisation. For example there is global workspace theory with its broadcast dynamics [undef, undefo], integrated information [undefz, undefu], higher-order representation [undefv], attention-schema modelling [undefq], predictive-processing and active-inference formalisms [undefp], and more general analyses of time and consciousness [undefr, undefaa]. The aim of this paper is not to adjudicate among theories, but to extract a common formal issue they face. Assume for the sake of argument that phenomenal content is modelled as a conjunction of grounded ingredients. If so, what does it then mean for that conjunction to be realised within the temporal extent of a subjective moment?
1.1 Results
  1. 1.
    Compositional grounding. I prove that any higher-layer statement can be grounded compositionally down to the base layer while preserving truth conditions (Theorem 1).
  2. 2.
    Subjective vs objective time. I formalise higher-layer ticks as maximal objective-time blocks of constant encoding, and show they can be arbitrarily sparse (Proposition 1).
  3. 3.
    Temporal lifting algebra. I define finite-window temporal lifting operators and prove the core non-commutation result underpinning the Temporal Gap (Theorem 3).
  4. 4.
    StrongSync vs WeakSync postulates. I state both postulates in terms of a higher-layer moment statement and its base-layer grounding (Definitions 19 and 20).
  5. 5.
    Concurrency capacity threshold. I show that architectures with limited simultaneous contribution can satisfy ingredient-wise occurrence while forbidding objective co-instantiation beyond capacity (Theorem 5).
  6. 6.
    A short argument supporting StrongSync. I briefly connect StrongSync to evidence linking consciousness to synchrony and effective connectivity (Section 7).
2 Stack Theory Primitives
I follow the core notation of Stack Theory [undefd, undeff].
Definition 1 (Environment and programs).
An environment is a nonempty set $\Phi$ of mutually exclusive states. A program is a set $p\subseteq\Phi$. A program $p$ is true at $\phi\in\Phi$ iff $\phi\in p$. Let $P:=2^{\Phi}$ be the set of all programs. A vocabulary is any finite $\mathfrak{v}\subseteq P$ (finite unless explicitly stated otherwise).
Intuition.
An environment is the set of possible states. A program is a set of states, each state representing a point in time and a smallest possible change in an environment, equating time with change [undeff]. A program is true at a state (because a state is an objective time) when the state is in that set. Conjunction is the intersection of the sets.
Definition 2 (Statements, truth sets, extensions).
The language generated by $\mathfrak{v}$ is
$\mathcal{L}_{\mathfrak{v}}:=\left\{\,l\subseteq\mathfrak{v}\;:\;l\ \text{is finite and }\bigcap_{p\in l}p\neq\emptyset\,\right\}.$
Elements $l\in\mathcal{L}_{\mathfrak{v}}$ are (conjunctive) statements. The truth set of $l$ is
$\mathrm{T}(l):=\bigcap_{p\in l}p\subseteq\Phi.$
I say $l$ is true at $\phi$ iff $\phi\in\mathrm{T}(l)$. By convention, $\mathrm{T}(\emptyset)=\Phi$.
A completion of $l$ is any $l^{\prime}\in\mathcal{L}_{\mathfrak{v}}$ with $l\subseteq l^{\prime}$. Define the extension of $l$ by $\mathrm{Ext}(l):=\{\,l^{\prime}\in\mathcal{L}_{\mathfrak{v}}:l\subseteq l^{\prime}\,\}$. For $X\subseteq\mathcal{L}_{\mathfrak{v}}$, write $\mathrm{Ext}(X):=\bigcup_{x\in X}\mathrm{Ext}(x)$.
Intuition.
A statement is a finite bundle of programs that must all be true at once. A completion adds more programs and makes the statement more specific. The extension is the collection of all completions that still include the original statement.
Definition 3 (Tasks and policies).
A task is a pair $\alpha=\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subseteq\mathcal{L}_{\mathfrak{v}}$ is a set of admissible input statements and $O_{\alpha}\subseteq\mathrm{Ext}(I_{\alpha})$ is a set of correct output statements. A policy is any statement $\pi\in\mathcal{L}_{\mathfrak{v}}$. A policy is correct for $\alpha$ iff
$\mathrm{Ext}(I_{\alpha})\cap\mathrm{Ext}(\pi)=O_{\alpha}.$
Let $\Pi_{\alpha}$ denote the set of correct policies for $\alpha$.
Intuition.
A task says which input statements are allowed and which output statements count as correct. A policy is correct when it allows exactly the correct outputs for the allowed inputs.
3 Grounding Stack Theory
Definition 4 (Abstractor).
Fix a base vocabulary $\mathfrak{v}\subseteq P$ and a policy $\pi\in\mathcal{L}_{\mathfrak{v}}$. The induced abstractor is the function
$\mathfrak{f}(\mathfrak{v},\pi):=\left\{\,q\in P\;:\;\exists o\in\mathrm{Ext}(\pi)\ \text{s.t.}\ \bigcap_{r\in o}r=q\,\right\}.$
Thus $\mathfrak{f}(\mathfrak{v},\pi)$ is a new vocabulary consisting of truth sets of completions compatible with $\pi$.
Intuition.
The abstractor takes completions of the policy and turns each completion into one set of states. The collection of these sets forms the vocabulary of the next layer. It’s basically a compression algorithm that takes completions and returns a set of equivalent programs.
Definition 5 (Stack).
A stack of depth $m$ is a sequence of uninstantiated task functions $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{m}\rangle$ such that for each $i<m$, $\lambda^{i+1}$ is a restriction of $\lambda^{i}$ (written $\lambda^{i+1}\sqsubset\lambda^{i}$). Formally, each $\lambda^{i}$ assigns to each vocabulary $\mathfrak{v}$ a task $\lambda^{i}(\mathfrak{v})$ as in Definition 3. The notation $\Pi_{\lambda^{i}(\mathfrak{v})}$ denotes the set of policies correct for that task. A stack state for such a stack is a pair of sequences
$(\pi^{0},\dots,\pi^{m-1})\quad\text{and}\quad(\mathfrak{v}^{0},\dots,\mathfrak{v}^{m})$
satisfying, for each $i<m$,
$\mathfrak{v}^{i+1}=\mathfrak{f}(\mathfrak{v}^{i},\pi^{i})\qquad\text{and}\qquad\pi^{i}\in\Pi_{\lambda^{i}(\mathfrak{v}^{i})}.$
Intuition.
At each layer you pick a policy that solves the task for that layer. That policy also defines the next vocabulary, which is the higher level way of describing what is going on.
3.1 Grounding a higher-layer into lower layers
Lemma 1 (Grounding through one abstractor).
Let $\mathfrak{v}\subseteq P$ and $\pi\in\mathcal{L}_{\mathfrak{v}}$, and set $\mathfrak{v}^{\prime}=\mathfrak{f}(\mathfrak{v},\pi)$. Then for every program $p\in\mathfrak{v}^{\prime}$ there exists a completion $o_{p}\in\mathrm{Ext}(\pi)$ such that $\mathrm{T}(o_{p})=p$. Moreover, for any statement $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$ there exists a statement $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$ with
$\mathrm{T}(l)=\mathrm{T}(\mathrm{ground}_{\pi}(l)).$
Proof.
By Definition 4, for each $p\in\mathfrak{v}^{\prime}$ there exists $o_{p}\in\mathrm{Ext}(\pi)$ with $\bigcap_{r\in o_{p}}r=p$, hence $\mathrm{T}(o_{p})=p$. Fix $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$. For each $p\in l$, choose such a witness completion $o_{p}$ and define
$\mathrm{ground}_{\pi}(l):=\bigcup_{p\in l}o_{p}\subseteq\mathfrak{v}.$
Then
$\mathrm{T}(\mathrm{ground}_{\pi}(l))$ $=\bigcap_{q\in\cup_{p\in l}o_{p}}q$
$=\bigcap_{p\in l}\bigcap_{q\in o_{p}}q$
$=\bigcap_{p\in l}\mathrm{T}(o_{p})$
$=\bigcap_{p\in l}p$
$=\mathrm{T}(l).$
Since $\mathrm{T}(l)\neq\emptyset$, also $\mathrm{T}(\mathrm{ground}_{\pi}(l))\neq\emptyset$. Hence $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$. ∎
Proof intuition.
Each macro-ingredient has a lower-layer completion with the same truth set. Replacing ingredients and intersecting preserves the overall truth conditions.
3.2 Compositional grounding
Definition 6 (Compositional grounding across a stack).
Fix a stack with vocabularies $\mathfrak{v}^{0},\dots,\mathfrak{v}^{m}$ and policies $\pi^{0},\dots,\pi^{m-1}$. For a statement $l^{m}\in\mathcal{L}_{\mathfrak{v}^{m}}$, define grounded statements $g^{m},\dots,g^{0}$ recursively by
$g^{m}$ $:=l^{m},$
$g^{i}$ $:=\mathrm{ground}_{\pi^{i}}(g^{i+1})\in\mathcal{L}_{\mathfrak{v}^{i}}\quad(i=m{-}1,\dots,0).$
Define the compositional grounding map $\mathrm{Ground}_{0\leftarrow m}$ by $\mathrm{Ground}_{0\leftarrow m}(l^{m}):=g^{0}$.
Intuition.
Grounding works like compiling a high level statement into lower level pieces. Replace each high level ingredient with a lower level statement that means the same thing. Repeating this down the stack gives a base level conjunction that is true in exactly the same cases.
Theorem 1 (Compositional grounding preserves truth conditions).
For any stack and any statement $l^{m}\in\mathcal{L}_{\mathfrak{v}^{m}}$,
$\mathrm{T}(\mathrm{Ground}_{0\leftarrow m}(l^{m}))=\mathrm{T}(l^{m}).$
Proof.
Apply Lemma 1 iteratively: $\mathrm{T}(g^{i})=\mathrm{T}(g^{i+1})$ for each $i<m$, hence $\mathrm{T}(g^{0})=\mathrm{T}(g^{m})=\mathrm{T}(l^{m})$. ∎
Proof intuition.
Iterate Lemma 1 down the stack and every grounding step preserves the truth set, so the composed grounding does too.
Remark 1 (Non-uniqueness (and why that is acceptable)).
Grounding is generally non-unique because the witness completions $o_{p}$ in Lemma 1 need not be unique. For Temporal Gap questions, what matters is that some base-layer grounded conjunction with the same truth set exists. Different groundings correspond to different decompositions of the same higher-layer content into base-level ingredients.
4 Stack-Time Semantics Module
This section is designed as a reusable addendum to Stack Theory. It defines induced layer time, window semantics, and temporal lifting operators in a way that can be cited by future work.
4.1 Objective time and layer time as a quotient
Definition 7 (Objective time and trajectories).
Objective time is represented by a discrete index $t\in\mathbb{N}$. An objective trajectory is a function $\tau:\mathbb{N}\to\Phi$ such that
$\tau(t{+}1)\neq\tau(t)\quad\text{for all }t\in\mathbb{N}.$
Intuition.
Objective time is a step counter for objective change. If nothing changes then there is no next objective index. Two events (statements) are simultaneous when they occur at the same objective index.
Definition 8 (Encoding map).
For a vocabulary $\mathfrak{v}\subseteq P$ and a state $\phi\in\Phi$, define
$\mathrm{Enc}_{\mathfrak{v}}(\phi):=\{\,p\in\mathfrak{v}:\phi\in p\,\}\subseteq\mathfrak{v}.$
Intuition.
Given a state and a vocabulary, the encoding collects all programs in that vocabulary that are true at that state. This gives the most detailed statement you can write in that vocabulary that still holds at that state.
Lemma 2 (Encoding yields the maximal true statement).
For any $\mathfrak{v}$ and $\phi\in\Phi$, $\mathrm{Enc}_{\mathfrak{v}}(\phi)\in\mathcal{L}_{\mathfrak{v}}$ and $\phi\in\mathrm{T}(\mathrm{Enc}_{\mathfrak{v}}(\phi))$. Moreover, if $l\in\mathcal{L}_{\mathfrak{v}}$ is true at $\phi$, then $l\subseteq\mathrm{Enc}_{\mathfrak{v}}(\phi)$.
Proof.
Since $\phi\in p$ for every $p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)$, I have $\phi\in\bigcap_{p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)}p=\mathrm{T}(\mathrm{Enc}_{\mathfrak{v}}(\phi))$, so the intersection is nonempty and $\mathrm{Enc}_{\mathfrak{v}}(\phi)\in\mathcal{L}_{\mathfrak{v}}$. If $l$ is true at $\phi$, then $\phi\in p$ for all $p\in l$, hence $p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)$ for all $p\in l$, so $l\subseteq\mathrm{Enc}_{\mathfrak{v}}(\phi)$. ∎
Proof intuition.
Every program in $\mathrm{Enc}_{\mathfrak{v}}(\phi)$ contains $\phi$, so their intersection contains $\phi$ and any statement true at $\phi$ can only use such programs.
Definition 9 (Induced layer trajectory and layer time).
Given an objective trajectory $\tau$ and a vocabulary $\mathfrak{v}$, define the induced layer trajectory
$\tau^{(\mathfrak{v})}(t):=\mathrm{Enc}_{\mathfrak{v}}(\tau(t))\in\mathcal{L}_{\mathfrak{v}}.$
Define the layer-tick map $\kappa_{\mathfrak{v}}:\mathbb{N}\to\mathbb{N}$ by $\kappa_{\mathfrak{v}}(0)=0$ and, for $t>0$,
$\kappa_{\mathfrak{v}}(t):=\begin{cases}\kappa_{\mathfrak{v}}(t{-}1)&\text{if }\tau^{(\mathfrak{v})}(t)=\tau^{(\mathfrak{v})}(t{-}1),\\ \kappa_{\mathfrak{v}}(t{-}1)+1&\text{otherwise.}\end{cases}$
Define an equivalence relation on objective times by $t\approx_{\mathfrak{v}}t^{\prime}$ iff $\kappa_{\mathfrak{v}}(t)=\kappa_{\mathfrak{v}}(t^{\prime})$. The corresponding layer time is the quotient set $\mathbb{N}/\!\approx_{\mathfrak{v}}$. Each equivalence class is a maximal contiguous block of objective times on which the induced macrostatement $\tau^{(\mathfrak{v})}$ is constant.
Intuition.
For each objective time step, write down the encoding at that step. If the encoding stays the same for several steps, treat those steps as one higher level moment. The tick counter increases only when the encoding changes, so higher level time advances only when something visible at that level changes.
Proposition 1 (Layer time can be arbitrarily sparse).
For any $N\in\mathbb{N}$ there exist $\Phi,\mathfrak{v},\tau$ such that the induced trajectory $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps. The objective-time gap between successive layer changes can be arbitrarily large.
Proof.
Let $\Phi=\{\phi_{t}:t\in\mathbb{N}\}$. Let $p:=\{\phi_{0},\phi_{1},\dots,\phi_{N}\}$ and take $\mathfrak{v}=\{p\}$. Define $\tau(t)=\phi_{t}$ for all $t\in\mathbb{N}$. Then $\tau$ is an objective trajectory because $\tau(t{+}1)\neq\tau(t)$ for all $t$. For all $t\leq N$ I have $\tau(t)\in p$, so $\tau^{(\mathfrak{v})}(t)=\{p\}$. I also have $\tau(N{+}1)\notin p$, so $\tau^{(\mathfrak{v})}(N{+}1)=\emptyset$. Hence $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps. ∎
Proof intuition.
Choose a vocabulary that cannot distinguish many distinct objective microstates. The induced macrostate then stays constant across many objective changes, so layer time barely advances.
4.2 Time and lifting operators
A subjective moment is commonly modelled as having nonzero temporal extent (the “specious present”) [undefr, undefaa]. To speak precisely about satisfaction within such a window, I introduce windowing maps and their induced temporal operators.
Definition 10 (Windowing maps).
A windowing map is a function $\mathrm{W}:\mathbb{N}\to\mathcal{P}(\mathbb{N})$ such that $\mathrm{W}(t)$ is finite and nonempty for every $t$. For horizon $\Delta\in\mathbb{N}$ and stride $s\in\mathbb{N}_{>0}$, the sliding window map is
$\mathrm{W}_{\Delta,s}(t)=\{st,st{+}1,\dots,st{+}\Delta\}.$
Intuition.
A windowing map picks, for each subjective index, a small set of objective time steps that belong to that moment. The parameter $\Delta$ controls how far the window extends, so the window contains $\Delta{+}1$ time steps. The parameter $s$ controls how far you shift the window each time.
Definition 11 (Ingredient-wise vs co-instantiated satisfaction).
Fix a windowing map $\mathrm{W}$, an objective trajectory $\tau$, and a subjective index $t$. For a program $p\subseteq\Phi$, define
$\mathrm{Occur}_{\mathrm{W}}(p,\tau,t)\iff\exists u\in\mathrm{W}(t)\ \big(\tau(u)\in p\big).$
For a statement $l\in\mathcal{L}_{\mathfrak{v}}$, define
$\mathrm{Occur}_{\mathrm{W}}(l,\tau,t)\iff\forall p\in l\ \mathrm{Occur}_{\mathrm{W}}(p,\tau,t).$
Define co-instantiation of the conjunction by
$\mathrm{CoInst}_{\mathrm{W}}(l,\tau,t)\iff\exists u\in\mathrm{W}(t)\ \big(\tau(u)\in\mathrm{T}(l)\big).$
Intuition.
Ingredient-wise means every ingredient becomes true at least once somewhere in the window. Co-instantiated means there is some time step in the window where all ingredients are true together.
Remark 2 (Quantifier structure).
For conjunctive $l$, ingredient-wise satisfaction is of the form $\forall p\in l\ \exists u\in\mathrm{W}(t)$, while co-instantiation is $\exists u\in\mathrm{W}(t)\ \forall p\in l$. The Temporal Gap is precisely the possible separation between these quantifier orders.
Definition 12 (Window environments and window trajectories).
Fix $\Delta\in\mathbb{N}$. Define the window environment $\Phi^{\Delta}:=\Phi^{\{0,1,\dots,\Delta\}}\cong\Phi^{\Delta+1}$ whose elements are windows $\sigma=(\phi_{0},\dots,\phi_{\Delta})$. Given stride $s$ and an objective trajectory $\tau$, define the induced window trajectory
$\tau^{\Delta,s}(t):=\big(\tau(st),\tau(st{+}1),\dots,\tau(st{+}\Delta)\big)\in\Phi^{\Delta}.$
Intuition.
Treat a whole window of $\Delta{+}1$ consecutive states as a single new state. This lets you test lifted statements one window at a time.
Definition 13 (Temporal lifts of programs).
For a program $p\subseteq\Phi$, define the lifted programs on $\Phi^{\Delta}$:
$\Diamond_{\Delta}p$ $=\{\,\sigma\in\Phi^{\Delta}:\exists k\leq\Delta\ \text{s.t.}\ \sigma(k)\in p\,\},$
$\Box_{\Delta}p$ $=\{\,\sigma\in\Phi^{\Delta}:\forall k\leq\Delta,\ \sigma(k)\in p\,\}.$
Intuition.
$\Diamond_{\Delta}p$ means $p$ is true at least once somewhere in the window. $\Box_{\Delta}p$ means $p$ is true at every time step in the window.
Definition 14 (Lifted statements).
For a statement $l\in\mathcal{L}_{\mathfrak{v}}$, define
$\Diamond_{\Delta}l$ $=\{\,\Diamond_{\Delta}p:p\in l\,\},$
$\Box_{\Delta}l$ $=\{\,\Box_{\Delta}p:p\in l\,\}.$
Then $\mathrm{T}(\Diamond_{\Delta}l)=\bigcap_{p\in l}\Diamond_{\Delta}p$ and similarly for $\Box$.
Intuition.
To lift a statement, lift each ingredient and then require all lifted ingredients to hold for the window. With $\Diamond_{\Delta}$ this means every ingredient shows up somewhere in the window. With $\Box_{\Delta}$ this means every ingredient stays true throughout the window.
Remark 3 (Internalising temporal semantics as an abstraction layer).
The window environment $\Phi^{\Delta}$ and the lifted vocabulary $\{\Diamond_{\Delta}p:p\in\mathfrak{v}\}$ (or $\{\Box_{\Delta}p\}$) together define a new abstraction layer in the native Stack-Theory sense. This is the formal bridge between “over time” constraints and pointwise Stack-Theory semantics.
4.3 The algebraic core
The following commutation and non-commutation facts are elementary, but spelling them out in Stack-Theory form turns the Temporal Gap into a concrete architectural question.
Theorem 2 (Universal lift commutes with conjunction).
For any statement $l\in\mathcal{L}_{\mathfrak{v}}$ and any $\Delta\in\mathbb{N}$,
$\Box_{\Delta}\mathrm{T}(l)\,=\,\mathrm{T}(\Box_{\Delta}l).$
Proof.
Fix any window $\sigma\in\Phi^{\Delta}$. Then
$\sigma\in\mathrm{T}(\Box_{\Delta}l)$ $\iff\forall p\in l,\ \sigma\in\Box_{\Delta}p$
$\iff\forall p\in l,\ \forall k\leq\Delta,\ \sigma(k)\in p$
$\iff\forall k\leq\Delta,\ \sigma(k)\in\bigcap_{p\in l}p$
$\iff\sigma\in\Box_{\Delta}\mathrm{T}(l).$
Proof intuition.
$\Box_{\Delta}$ is “for all positions in the window”. Since $\forall$ distributes over conjunction, $\Box_{\Delta}$ commutes with $\wedge$.
Theorem 3 (Existential lift does not commute with conjunction).
For any statement $l\in\mathcal{L}_{\mathfrak{v}}$ and any $\Delta\in\mathbb{N}$,
$\Diamond_{\Delta}\mathrm{T}(l)\;\subseteq\;\mathrm{T}(\Diamond_{\Delta}l).$
Moreover, for $\Delta\geq 1$ there exist environments and statements for which the inclusion is strict.
Proof.
If $\sigma\in\Diamond_{\Delta}\mathrm{T}(l)$ then for some $k$ I have $\sigma(k)\in\mathrm{T}(l)=\bigcap_{p\in l}p$. Hence $\sigma\in\Diamond_{\Delta}p$ for every $p\in l$ and therefore $\sigma\in\bigcap_{p\in l}\Diamond_{\Delta}p=\mathrm{T}(\Diamond_{\Delta}l)$.
For strictness when $\Delta\geq 1$, let $\Phi=\{a,b,c\}$ and $\mathfrak{v}=\{p,q\}$ with $p=\{a,c\}$ and $q=\{b,c\}$. Then $l=\{p,q\}\in\mathcal{L}_{\mathfrak{v}}$ because $p\cap q=\{c\}\neq\emptyset$. Take $\Delta=1$ and window $\sigma=(a,b)$. Then $\sigma\in\Diamond_{1}p$ and $\sigma\in\Diamond_{1}q$, so $\sigma\in\mathrm{T}(\Diamond_{1}l)$. But $\sigma\notin\Diamond_{1}\mathrm{T}(l)=\Diamond_{1}(p\cap q)$ because neither $a$ nor $b$ lies in $p\cap q=\{c\}$. ∎
Proof intuition.
$\Diamond_{\Delta}$ is “there exists a position”. Since $\exists$ does not distribute over conjunction, ingredients may occur at different positions without co-occurring.
Remark 4 (Degenerate equality cases).
If $\Delta=0$ (a window is a single instant) or $\lvert l\rvert\leq 1$ (no nontrivial conjunction), then
$\Diamond_{\Delta}\mathrm{T}(l)\,=\,\mathrm{T}(\Diamond_{\Delta}l).$
Thus strict separation requires $\Delta\geq 1$ and at least two ingredients.
Remark 5 (Reading the theorem back into window semantics).
In the sliding-window case, Definition 11 corresponds to
$\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(l,\tau,t)$ $\iff\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}l),$
$\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(l,\tau,t)$ $\iff\tau^{\Delta,s}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
Thus Theorem 3 is exactly the claim that ingredient-wise satisfaction is strictly weaker than co-instantiation in general.
4.4 When does $\Diamond$ commute with conjunction?
Definition 15 (Within-window persistence along a trajectory).
Fix $\Delta\in\mathbb{N}$ and an objective trajectory $\tau:\mathbb{N}\to\Phi$. A program $p\subseteq\Phi$ is $(\tau,\Delta)$-persistent if for all $t\in\mathbb{N}$ and all indices $k\leq k^{\prime}\leq\Delta$,
$\tau(t+k)\in p\ \Longrightarrow\ \tau(t+k^{\prime})\in p.$
Intuition.
Along the trajectory, if $p$ becomes true at some point inside a $\Delta$ window, then $p$ stays true for the rest of that same window.
Proposition 2 (A sufficient condition for co-instantiation along a trajectory).
Fix $\Delta\in\mathbb{N}$ and an objective trajectory $\tau$. Let $l\in\mathcal{L}_{\mathfrak{v}}$ be a statement such that every ingredient program $p\in l$ is $(\tau,\Delta)$-persistent. Then for all $t$,
$\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l)\quad\Longrightarrow\quad\tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
Proof.
Assume $\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l)$. Then for each $p\in l$ there exists $k_{p}\leq\Delta$ with $\tau(t+k_{p})\in p$. Let $k^{\ast}:=\max_{p\in l}k_{p}$. By $(\tau,\Delta)$-persistence of each $p$, from $\tau(t+k_{p})\in p$ and $k_{p}\leq k^{\ast}\leq\Delta$ I obtain $\tau(t+k^{\ast})\in p$ for all $p\in l$. Hence $\tau(t+k^{\ast})\in\bigcap_{p\in l}p=\mathrm{T}(l)$, so $\tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(l)$. ∎
Proof intuition.
With persistence, once each ingredient appears it remains true. Choosing the latest appearance time yields a single instant where all ingredients overlap.
4.5 Finite-horizon temporal semantics
Finite-horizon (history-dependent) temporal semantics can itself be treated as an abstraction layer. A horizon-$\Delta$ pathwise abstractor can be compiled into a pointwise abstractor on an augmented state, so Stack-Theory tools apply without modifying the underlying semantics.
Definition 16 (Pointwise and horizon-$\Delta$ pathwise abstractors).
Let $\Psi$ be any higher-level state space. A pointwise abstractor is a function $A:\Phi\to\Psi$. A pathwise abstractor of horizon $\Delta$ is a function $A^{(\Delta)}:\Phi^{\Delta}\to\Psi$.
Intuition.
A pointwise abstractor looks only at the current state. A horizon $\Delta$ pathwise abstractor looks at the recent $\Delta{+}1$ states, so it can use short term history.
Theorem 4 (History-state (state augmentation) compilation).
Fix $\Delta\in\mathbb{N}$ and let $A^{(\Delta)}:\Phi^{\Delta}\to\Psi$ be a pathwise abstractor. Define the augmented environment $\widetilde{\Phi}:=\Phi^{\Delta}$ and the pointwise abstractor $\widetilde{A}:\widetilde{\Phi}\to\Psi$ by $\widetilde{A}(\sigma):=A^{(\Delta)}(\sigma)$. Then for every objective trajectory $\tau:\mathbb{N}\to\Phi$ and every $t$, with augmented trajectory $\widetilde{\tau}(t):=\tau^{\Delta,1}(t)\in\widetilde{\Phi}$ I have
$\widetilde{A}(\widetilde{\tau}(t))=A^{(\Delta)}(\tau^{\Delta,1}(t)).$
Proof.
Immediate from the definitions of $\widetilde{A}$ and $\widetilde{\tau}$. ∎
Proof intuition.
Encode the last $\Delta$ microstates into an augmented state so any horizon-$\Delta$ (history-dependent) abstractor becomes pointwise on the augmented state.
5 StrongSync and WeakSync Postulates
The above algebra does not decide which window semantics corresponds to phenomenality. It clarifies what must be decided. I’ll now state StrongSync and WeakSync in terms of a higher-layer moment statement and its base-layer grounding across a stack state.
Assumption 1 (Subjective moments live at some layer).
Fix a stack state with vocabularies $\mathfrak{v}^{0},\dots,\mathfrak{v}^{m}$. Assume a candidate subjective moment is represented by a conjunctive statement $l^{m}\in\mathcal{L}_{\mathfrak{v}^{m}}$.
Definition 17 (Phenomenal realisation predicate).
Let $\mathrm{PhenReal}(l^{m},\tau,t)$ mean “moment statement $l^{m}$ is phenomenally realised at (higher-layer) time $t$ along objective trajectory $\tau$.” I treat $\mathrm{PhenReal}$ as primitive, constrained by StrongSync/WeakSync.
Intuition.
$\mathrm{PhenReal}$ names the claim that moment statement $l^{m}$ is experienced at subjective time $t$ along the run $\tau$. Here $l^{m}$ can include first, second, and third order selves and causal identities for objects and properties [undeff].
Definition 18 (Base grounding of a moment statement).
Let $g^{0}:=\mathrm{Ground}_{0\leftarrow m}(l^{m})$ be the base-layer grounding of $l^{m}$ (Definition 6). By Theorem 1, $\mathrm{T}(g^{0})=\mathrm{T}(l^{m})$. I call $g^{0}$ the grounded conjunction of the moment statement.
Intuition.
$g^{0}$ is the low level conjunction that corresponds to the higher level moment statement $l^{m}$. It is built by expanding $l^{m}$ down to the base layer. It is true in exactly the same situations as $l^{m}$.
Definition 19 (StrongSync (objective co-instantiation required)).
Fix a windowing map for the phenomenal layer, typically the sliding window $\mathrm{W}_{\Delta,s}$ of horizon $\Delta$ and stride $s$ (Section 4.2). StrongSync postulates that for all $l^{m},\tau,t$,
$\mathrm{PhenReal}(l^{m},\tau,t)$ $\Rightarrow\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)$
$\text{(equivalently }\tau^{\Delta,s}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\text{).}$
If the moment is experienced, then within the corresponding subjective window there exists an objective time-slice at which the entire grounded conjunction holds.
Intuition.
StrongSync says that if the moment is experienced, then within the window there is at least one objective time step where all grounded ingredients are true together.
Definition 20 (WeakSync (objective smearing permitted)).
Fix the same windowing map $\mathrm{W}_{\Delta,s}$ as in Definition 19. WeakSync postulates the following.
  1. 1.
    Ingredient-wise necessity. For all $l^{m},\tau,t$,
    $\mathrm{PhenReal}(l^{m},\tau,t)\Rightarrow\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)$
    $\text{(equivalently}\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\text{).}$
  2. 2.
    Smearing permitted. There exist $l^{m},\tau,t$ such that
    $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)\wedge\neg\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t),$
    equivalently $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\wedge\tau^{\Delta,s}(t)\notin\Diamond_{\Delta}\mathrm{T}(g^{0})$.
In plain English this allows a moment to be experienced without any instant realising the full grounded conjunction.
Intuition.
WeakSync says that if the moment is experienced, then each grounded ingredient must occur somewhere in the window. It allows that there may be no single objective time step where they are all true together.
6 Architectural Consequences
A polycomputer simultaneously executes multiple functions using the same parts or physical substrate [undefk]. A liquid brain such as an ant colony does not rely on a persistent structure to compute, but instead computes through movement [undefx]. In contrast, a “solid brain” is one that has a persistent structure that can support, for example, a bio-electric abstraction layer with synchronous broadcast [undeff] (like a network of telegraph cables). I now translate “single-thread sequential emulation” versus “synchronous, solid-brained, polycomputational multi-contributor realisation” into a minimal architectural invariant with a bound on simultaneous causal contribution. The Temporal Gap determines whether that bound is consciousness-relevant.
6.1 Concurrency capacity as an invariant
Definition 21 (Contributor model).
Fix $n\geq 2$ potential contributors and write $n]:=\{1,\dots,n\}$. Define a base environment of microstates
$\Phi:=X\times\mathcal{P}([n]),$
where $x\in X$ is a “data” component and $A\subseteq[n]$ is the set of currently active contributors at that objective instant. For each $i\in[n]$ define the program
$p_{i}:=\{(x,A)\in\Phi:i\in A\}.$
Let $l_{n}:=\{p_{1},\dots,p_{n}\}$ and note $\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$.
Intuition.
The contributor model records which contributors are active at each time step. A conjunction that mentions many contributors can only be true when all of those contributors are active at the same time step.
Definition 22 (Concurrency capacity).
An architecture has concurrency capacity $c\in\{1,\dots,n\}$ if its admissible trajectories $\tau:\mathbb{N}\to\Phi$ satisfy (writing $\tau(t)=(x_{t},A_{t})$ for the data and contributor-set components)
$\lvert A_{t}\rvert\leq c\quad\text{for all objective times }t.$
The case $c=1$ formalises a sequential single-thread update model (one active contributor per tick). The case $c\geq n$ formalises synchronous polycomputing at the relevant scale (all contributors may be active together).
Intuition.
Capacity $c$ is the maximum number of contributors that can be active at one time step. If the grounded content needs more than $c$ contributors active together, then StrongSync cannot hold on that architecture.
Theorem 5 (Synchronous threshold for co-instantiation).
Fix $n\geq 2$ and content statement $l_{n}$. If an architecture has concurrency capacity $c<n$, then along every admissible trajectory $\tau$ and for every horizon $\Delta$ I have
$\tau^{\Delta,1}(t)\notin\Diamond_{\Delta}\mathrm{T}(l_{n})\quad\text{for all }t.$
However, if the architecture is such that every contributor is activated at least once within some horizon-$\Delta$ window (i.e. for some $t$ and all $i$ there exists $k_{i}\leq\Delta$ with $i\in A_{t+k_{i}}$), then
$\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l_{n}).$
Thus the architecture admits ingredient-wise satisfaction but forbids objective co-instantiation.
Proof.
If $c<n$, then at any time $t$ I have $\lvert A_{t}\rvert\leq c<n$, so $A_{t}\neq[n]$. Hence $\tau(t)\notin\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$ for all $t$. Therefore no window contains an element of $\mathrm{T}(l_{n})$, so $\Diamond_{\Delta}\mathrm{T}(l_{n})$ is never satisfied. For the second claim, if each $i$ appears in $A_{t+k_{i}}$ for some $k_{i}\leq\Delta$, then the window $\sigma=\tau^{\Delta,1}(t)$ satisfies $\sigma\in\Diamond_{\Delta}p_{i}$ for all $i$, hence $\sigma\in\bigcap_{i}\Diamond_{\Delta}p_{i}=\mathrm{T}(\Diamond_{\Delta}l_{n})$. ∎
Proof intuition.
If at most $c$ contributors can be active, an $n$-way conjunction with $n>c$ is never true at an instant, though each ingredient can appear somewhere in a window.
Corollary 1 (Single-thread CPU vs synchronous polycomputer (formal)).
Let a moment statement $l^{m}$ have base grounding $g^{0}$ whose ingredients include a conjunctive substatement equivalent to $l_{n}$ for some $n\geq 2$. Fix the phenomenal windowing map $\mathrm{W}_{\Delta,s}$ used in StrongSync/WeakSync.
  1. 1.
    Under StrongSync, any sequential-update architecture with capacity $c<n$ cannot phenomenally realise $l^{m}$, because $\Diamond_{\Delta}\mathrm{T}(g^{0})$ fails for the relevant grounded substatement (Theorem 5).
  2. 2.
    Under WeakSync, the same sequential architecture remains a candidate in principle, because $\mathrm{T}(\Diamond_{\Delta}g^{0})$ may hold without $\Diamond_{\Delta}\mathrm{T}(g^{0})$.
  3. 3.
    Any synchronous architecture with capacity $c\geq n$ can satisfy $\Diamond_{\Delta}\mathrm{T}(g^{0})$ whenever it activates the required contributors simultaneously, hence can host the moment under both StrongSync and WeakSync.
Remark 6 (Connection to valence-first Stack Theory).
My valence-first application of Stack Theory treats certain phenomenal configurations as synchronised, causally efficacious conjunctions over many ingredients [undeff]. The capacity model above isolates one minimal implementation parameter, simultaneous contribution, that controls whether such conjunctions can be instantiated at any objective instant versus only “smeared” across a window.
7 Evidence for StrongSync
StrongSync and WeakSync predict different empirical signatures within a short subjective window. Either consciousness requires a window-local episode of coordinated co-engagement (objective co-instantiation), or it can tolerate mere ingredient-wise activation distributed across the window. Several classic paradigms are more naturally described by the coordination view. In masking, conscious perception has been reported to correlate with transient long-range gamma phase synchrony across cortical areas, even when local activity changes are similar across seen versus unseen stimuli [undeft]. In non-REM sleep, TMS-EEG responses can become strong yet local, failing to propagate, consistent with a breakdown of effective connectivity when consciousness fades [undefs]. More generally, perturbational measures such as PCI track level of consciousness across wake/sleep/anesthesia and disorders of consciousness, consistent with a requirement for integrated, differentiated system-level dynamics [undefm]. Taken together, these findings do not prove StrongSync, but they do make the existential permissiveness of WeakSync much less compelling. Conscious level appears to covary with the presence (and disappearance with the collapse) of a temporally coordinated, system-level regime. Hence when I draw architectural conclusions below, I treat StrongSync as the default working hypothesis. This step remains defeasible.
8 Attribution and Implications
The machine consciousness context puts special weight on attribution/measurement under uncertainty and implementability constraints. The Temporal Gap formalism makes a concrete measurement target of the gap between ingredient coverage and co-instantiation for a grounded conjunction.
8.1 Instrumentable separation metrics
Fix a grounded conjunction $g^{0}$ and an objective trajectory $\tau$ (possibly generated under controlled conditions). Define the minimal horizon needed for ingredient coverage versus co-instantiation:
$w_{\mathrm{ing}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\},$
$w_{\mathrm{co}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\}.$
I use the convention $\inf\emptyset=\infty$ (so these live in $\mathbb{N}\cup\{\infty\}$). The strict Temporal Gap pattern is $w_{\mathrm{ing}}<w_{\mathrm{co}}$ (or $w_{\mathrm{co}}=\infty$).
These are, in principle, instrumentable when the base-layer ingredients are defined in terms of architectural state variables (e.g. active contributor sets). In the broader context of AI consciousness assessment, this kind of architectural metric aligns with the view that behaviour alone may be insufficient and that concrete indicator properties should be checked against candidate theories [undefl].
8.2 Contrast-class experiment
Construct paired systems that are behaviourally matched at a reporting layer (same task performance and report interface) but differ in base-layer concurrency capacity:
  1. 1.
    Synchronous substrate: implement a bounded-latency global tick and allow high concurrency at the relevant scale (capacity $c\geq n$).
  2. 2.
    Sequential substrate: implement the same macro update rule but enforce small $c$ by serialising contributor updates.
Then select candidate contents $l^{m}$ whose groundings include large conjunctions (Section 6), and measure $w_{\mathrm{ing}}$ and $w_{\mathrm{co}}$. StrongSync predicts that phenomenality (if present) tracks bounded $w_{\mathrm{co}}$ at the relevant grounding layer. WeakSync predicts it may track $w_{\mathrm{ing}}$ alone.
8.3 Normative significance under uncertainty
If StrongSync is even plausible, then behavioural equivalence between a synchronous system and a sequential emulation does not settle whether they are phenomenally equivalent. This creates an ethically salient zone of reasonable disagreement about moral status, particularly when a system could have valenced experience. A precautionary stance toward sentience under uncertainty has been defended in nearby contexts [undefi]. Within my framework, one concrete precautionary recommendation is to treat large Temporal-Gap regimes ($w_{\mathrm{co}}\gg w_{\mathrm{ing}}$ or $w_{\mathrm{co}}=\infty$) as a high-uncertainty region for attribution, motivating conservative deployment and experimentation policies until the StrongSync/WeakSync question is better resolved.
9 Conclusion
I provided a Stack-Theory-compatible Stack-Time Semantics addendum and used it to sharpen the Temporal Gap into a precise algebraic and architectural question. Compositional grounding (Theorem 1) ensures that any higher-layer phenomenal statement has a base-layer grounded conjunction with identical truth conditions. Induced layer time partitions objective time into maximal constant macro-blocks (a change-point quotient) and can be arbitrarily sparse (Proposition 1), so subjective moments need not correspond to single objective instants. Against that backdrop, the Temporal Gap becomes the non-commutation pattern
$\Diamond_{\Delta}\mathrm{T}(l)\subseteq\mathrm{T}(\Diamond_{\Delta}l),$
with strict inclusion in general for $\Delta\geq 1$ and multi-ingredient statements (Theorem 3, Remark 4). This yields a clean separation between ingredient-wise window occurrence and objective co-instantiation.
StrongSync and WeakSync are then competing postulates about which notion constrains phenomenality. Section 7 notes that conscious level often covaries with temporally coordinated, system-level dynamics (synchrony/connectivity/complexity), which tentatively favours StrongSync over WeakSync.
Finally, I proved that a simple architectural invariant (concurrency capacity) can force the strict Temporal Gap pattern for large conjunctions (Theorem 5). If StrongSync is correct and if a phenomenal content’s base grounding requires a conjunction larger than a substrate’s simultaneous causal contribution can support, then that content cannot be phenomenally realised on that substrate. Under StrongSync, software consciousness on a strictly sequential (capacity-$1$) substrate is impossible for any content whose grounding requires two or more simultaneous contributors. Real machines are not perfectly sequential at the physical level. Thus the practical question becomes which physical concurrency scale is relevant for grounding. Consciousness according to Stack Theory requires a “tapestry of valence”, supported by a solid-brained polycomputing architecture, in which each co-instantiated ingredient simultaneously acts upon the others as they act on it, all linked and synchronised. An extreme concurrency capacity may be needed, and still concurrency is only one (important) aspect of this. The central conclusion remains that consciousness attribution is hardware-sensitive, and surface-level observable behavioural equivalence alone cannot wash out the Temporal Gap.

### v2

@@
 Whether machines can be conscious depends not only on what they compute, but when they compute it.
-Most deployed artificial systems realise their functions via sequential or time-multiplexed updates.
-Conscious experience appears unified and simultaneous.
-I show that this difference matters formally.
+Most deployed artificial systems realise their functions via sequential or time-multiplexed updates, yet a moment of conscious experience feels unified and simultaneous.
+I prove that this difference matters.
 I augment Stack Theory with algebraic laws relating within time-window constraint satisfaction to conjunction.
-I introduce a precise temporal semantics over windowed trajectories $\tau^{\Delta,s}$ and prove that existential temporal realisation $\Diamond_{\Delta}$ does not preserve conjunction.
+I introduce a temporal semantics over windowed trajectories $\tau_{\Delta}$ and prove that existential temporal realisation $\Diamond_{\Delta}$ does not preserve conjunction.
 A system can realise all the ingredients of experience across time without ever instantiating the experienced conjunction itself.
-I then distinguish two postulates.
-StrongSync requires objective co-instantiation of the grounded conjunction within the window, while WeakSync permits temporal “smearing”.
-I formalise concurrency-capacity to measure what is needed to satisfy StrongSync.
-Finally, I review neurophysiological evidence suggesting that consciousness depends on phase synchrony and effective connectivity, and that loss of consciousness is often associated with its breakdown.
-This evidence makes WeakSync less plausible.
-Under StrongSync, software consciousness on strictly sequential substrates is impossible for contents whose grounding requires two or more simultaneous contributors.
-The more parts from which simultaneous contribution required, the more concurrency capacity is required.
+I then distinguish two postulates, Chord and Arpeggio.
+Chord is the position that conscious unity requires objective co-instantiation of the grounded conjunction within the window, like a musical chord.
+Arpeggio only needs the ingredients to occur within window, like a melody.
+I formalise concurrency-capacity to measure what is needed to satisfy co-instantiation.
+Finally, I review neurophysiological evidence suggesting that consciousness depends on phase synchrony and effective connectivity, and that loss of consciousness is associated with its breakdown.
+Under Chord, software consciousness on strictly sequential substrates is impossible for contents whose grounding requires two or more simultaneous contributors.
 The hardware matters.
-Consciousness attribution therefore requires architectural inspection, not just functional performance.
 1 Introduction
 A central open question in machine consciousness is whether behaviourally and functionally matched systems must be matched in phenomenal realisation.
-Stack Theory seeks to approach this question by formalising cognitive processes in terms of abstraction layers [undeff, undefb, undefh].
-Various aspects of incumbent theories of consciousness are formalised [undefw, undefj, undefn, undefv, undefy, undef], organised into a hierarchy of causal identities [undefb], and derived as a necessary consequence of a generalisation-optimal learning theory that makes valence primary [undefe, undefg, undefc, undefa].
+Stack Theory approaches this by formalising cognitive processes in terms of abstraction layers [7, 4, 2].
+Various aspects of incumbent theories of consciousness are formalised [24, 12, 16, 23, 26, 1], organised into a hierarchy of causal identities [4], and derived as a necessary consequence of a generalisation-optimal learning theory that makes valence primary [6, 8, 5, 3].
 Those necessary ingredients of consciousness are beyond the scope of this paper.
 Here I am concerned with a question of sufficiency.
-Crucially, Stack Theory shows how the behaviour of two systems can appear identical at one level of abstraction, while being different in the levels below.
-This means that conscious experience can be purely functional, but still depend on more than surface level function.
-It can depend not just on what is computed, but when and how.
-The ingredients of what looks like a single experienced moment at a high level of abstraction may be smeared across objective micro-time at a lower level of abstraction, such that no objective time-slice contains the full grounded conjunction (meaning the parts of a subjective moment of consciousness exist at different objective times, never together).
-If the ingredients of a subjective moment need not synchronise in objective time for that moment to be experienced, then current hardware is sufficient to host machine consciousness.
-That also means a “liquid-brained” swarm intelligence [undefx] like an ant colony or a population of humans can be conscious.
-However, if synchronised broadcast and causal co-dependence matter, then far fewer systems can be conscious.
-The hardware matters, because the parts of a subjective conscious moment need to be connected in objective time to be experienced.
-This question of architectural sufficiency is called The Temporal Gap.
-The Temporal Gap was proposed in earlier work [undeff], but was not rigorously formalised.
-Here I seek to rectify that oversight with two kinds of contribution.
-Formally, I add a Stack-Time Semantics module to Stack Theory.
-This module includes induced layer time, window environments, temporal lifting operators, and algebraic laws relating “within-window” satisfaction to conjunction.
+Stack Theory shows how the behaviour of two systems can appear identical at one level of abstraction, while being different in the levels below.
+Conscious experience can be purely functional, but still depend on more than surface level function—not just on what is computed, but when and how.
+The ingredients of what looks like a single experienced moment at a high level of abstraction may be smeared across objective micro-time at a lower level, such that no objective time-slice contains the full grounded conjunction.
+If the ingredients of a subjective moment need not synchronise in objective time, then current hardware suffices to host machine consciousness—and a “liquid-brained” swarm intelligence [25] like an ant colony can be conscious.
+But if synchronised broadcast and causal co-dependence matter, then far fewer systems qualify.
+This question of architectural sufficiency is the Temporal Gap.
+The Temporal Gap was proposed in earlier work [7], but was not rigorously formalised.
+Here I add a Stack-Time Semantics module to Stack Theory, including induced layer time, window environments, temporal lifting operators, and algebraic laws relating “within-window” satisfaction to conjunction.
 Using this, I refine the Temporal Gap into a testable architectural question about synchrony and concurrency.
 Relation to existing consciousness theories.
-Beyond Stack Theory, there are many scientific theories which emphasise some form of integration or unity across time and causal organisation.
-For example there is global workspace theory with its broadcast dynamics [undef, undefo], integrated information [undefz, undefu], higher-order representation [undefv], attention-schema modelling [undefq], predictive-processing and active-inference formalisms [undefp], and more general analyses of time and consciousness [undefr, undefaa].
-The aim of this paper is not to adjudicate among theories, but to extract a common formal issue they face.
-Assume for the sake of argument that phenomenal content is modelled as a conjunction of grounded ingredients.
-If so, what does it then mean for that conjunction to be realised within the temporal extent of a subjective moment?
+Beyond Stack Theory, many scientific theories emphasise some form of integration or unity across time and causal organisation: global workspace theory [1, 17], integrated information [27, 22], higher-order representation [23], attention-schema modelling [29], predictive-processing and active-inference [18], and more general analyses of time and consciousness [19, 28].
+The aim here is not to adjudicate among theories, but to extract a common formal issue: if phenomenal content is modelled as a conjunction of grounded ingredients, what does it mean for that conjunction to be realised within the temporal extent of a subjective moment?
 1.1 Results
   1. 1.
-    Compositional grounding. I prove that any higher-layer statement can be grounded compositionally down to the base layer while preserving truth conditions (Theorem 1).
+    Compositional grounding. Any higher-layer statement can be grounded compositionally to the base layer preserving truth conditions (Theorem 1).
   2. 2.
-    Subjective vs objective time. I formalise higher-layer ticks as maximal objective-time blocks of constant encoding, and show they can be arbitrarily sparse (Proposition 1).
+    Subjective vs objective time. Higher-layer ticks are maximal objective-time blocks of constant encoding, and can be arbitrarily sparse (Proposition 1).
   3. 3.
-    Temporal lifting algebra. I define finite-window temporal lifting operators and prove the core non-commutation result underpinning the Temporal Gap (Theorem 3).
+    Temporal lifting algebra. The core non-commutation result underpinning the Temporal Gap (Theorem 3).
   4. 4.
-    StrongSync vs WeakSync postulates. I state both postulates in terms of a higher-layer moment statement and its base-layer grounding (Definitions 19 and 20).
+    Chord vs Arpeggio postulates. Both postulates stated in terms of a higher-layer moment statement and its base-layer grounding (Definitions 16 and 17).
   5. 5.
-    Concurrency capacity threshold.
-I show that architectures with limited simultaneous contribution can satisfy ingredient-wise occurrence while forbidding objective co-instantiation beyond capacity (Theorem 5).
+    Concurrency capacity threshold. Architectures with limited simultaneous contribution can satisfy ingredient-wise occurrence while forbidding co-instantiation beyond capacity (Theorem 4).
   6. 6.
-    A short argument supporting StrongSync. I briefly connect StrongSync to evidence linking consciousness to synchrony and effective connectivity (Section 7).
+    Evidence supporting Chord. Co-instantiation requirements are connected to evidence linking consciousness to synchrony and effective connectivity (Section 7).
 2 Stack Theory Primitives
-I follow the core notation of Stack Theory [undefd, undeff].
+I follow the core notation of Stack Theory [9, 7].
 Definition 1 (Environment and programs).
@@
 Let $P:=2^{\Phi}$ be the set of all programs.
-A vocabulary is any finite $\mathfrak{v}\subseteq P$ (finite unless explicitly stated otherwise).
+A vocabulary is any finite $\mathfrak{v}\subseteq P$.
 Intuition.
 An environment is the set of possible states.
-A program is a set of states, each state representing a point in time and a smallest possible change in an environment, equating time with change [undeff].
-A program is true at a state (because a state is an objective time) when the state is in that set.
-Conjunction is the intersection of the sets.
+A program is a set of states, each state representing a point in time and a smallest possible change in an environment, equating time with change [7].
+A program is true at a state when the state is in that set.
+Conjunction is set intersection.
 Definition 2 (Statements, truth sets, extensions).
@@
 A statement is a finite bundle of programs that must all be true at once.
-A completion adds more programs and makes the statement more specific.
-The extension is the collection of all completions that still include the original statement.
+A completion adds more programs, making the statement more specific.
+The extension collects all completions that still include the original statement.
 Definition 3 (Tasks and policies).
@@
 Intuition.
-A task says which input statements are allowed and which output statements count as correct. A policy is correct when it allows exactly the correct outputs for the allowed inputs.
+A task specifies which inputs are allowed and which outputs count as correct. A policy is correct when it selects exactly the correct outputs for the allowed inputs.
 3 Grounding Stack Theory
@@
 Intuition.
-The abstractor takes completions of the policy and turns each completion into one set of states.
-The collection of these sets forms the vocabulary of the next layer.
-It’s basically a compression algorithm that takes completions and returns a set of equivalent programs.
+The abstractor takes completions of the policy and turns each into one set of states.
+The collection of these sets forms the vocabulary of the next layer—a compression that maps completions to equivalent programs.
 Definition 5 (Stack).
 A stack of depth $m$ is a sequence of uninstantiated task functions $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{m}\rangle$ such that for each $i<m$, $\lambda^{i+1}$ is a restriction of $\lambda^{i}$ (written $\lambda^{i+1}\sqsubset\lambda^{i}$).
-Formally, each $\lambda^{i}$ assigns to each vocabulary $\mathfrak{v}$ a task $\lambda^{i}(\mathfrak{v})$ as in Definition 3.
-The notation $\Pi_{\lambda^{i}(\mathfrak{v})}$ denotes the set of policies correct for that task.
-A stack state for such a stack is a pair of sequences
+Each $\lambda^{i}$ assigns to each vocabulary $\mathfrak{v}$ a task $\lambda^{i}(\mathfrak{v})$.
+The notation $\Pi_{\lambda^{i}(\mathfrak{v})}$ denotes the set of correct policies for that task.
+A stack state is a pair of sequences
 $(\pi^{0},\dots,\pi^{m-1})\quad\text{and}\quad(\mathfrak{v}^{0},\dots,\mathfrak{v}^{m})$
@@
 Intuition.
-At each layer you pick a policy that solves the task for that layer. That policy also defines the next vocabulary, which is the higher level way of describing what is going on.
+At each layer you pick a policy that solves the task for that layer. That policy also defines the next vocabulary, the higher level description of what is going on.
 3.1 Grounding a Higher-Layer into Lower Layers
@@
 Then for every program $p\in\mathfrak{v}^{\prime}$ there exists a completion $o_{p}\in\mathrm{Ext}(\pi)$ such that $\mathrm{T}(o_{p})=p$.
-Moreover, for any statement $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$ there exists a statement $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$ with
-$\mathrm{T}(l)=\mathrm{T}(\mathrm{ground}_{\pi}(l)).$
-Proof.
-By Definition 4, for each $p\in\mathfrak{v}^{\prime}$ there exists $o_{p}\in\mathrm{Ext}(\pi)$ with $\bigcap_{r\in o_{p}}r=p$, hence $\mathrm{T}(o_{p})=p$.
-Fix $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$.
-For each $p\in l$, choose such a witness completion $o_{p}$ and define
-$\mathrm{ground}_{\pi}(l):=\bigcup_{p\in l}o_{p}\subseteq\mathfrak{v}.$
-Then
-$\mathrm{T}(\mathrm{ground}_{\pi}(l))$ $=\bigcap_{q\in\cup_{p\in l}o_{p}}q$
-$=\bigcap_{p\in l}\bigcap_{q\in o_{p}}q$
-$=\bigcap_{p\in l}\mathrm{T}(o_{p})$
-$=\bigcap_{p\in l}p$
-$=\mathrm{T}(l).$
-Since $\mathrm{T}(l)\neq\emptyset$, also $\mathrm{T}(\mathrm{ground}_{\pi}(l))\neq\emptyset$. Hence $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$. ∎
-Proof intuition.
-Each macro-ingredient has a lower-layer completion with the same truth set. Replacing ingredients and intersecting preserves the overall truth conditions.
+Moreover, for any statement $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$ there exists a statement $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$ with $\mathrm{T}(l)=\mathrm{T}(\mathrm{ground}_{\pi}(l))$.
+Proof. See Supplementary Material. Each macro-ingredient has a lower-layer completion with the same truth set; replacing ingredients and intersecting preserves truth conditions. ∎
 3.2 Compositional Grounding
@@
 Replace each high level ingredient with a lower level statement that means the same thing.
-Repeating this down the stack gives a base level conjunction that is true in exactly the same cases.
+Repeating this down the stack gives a base level conjunction true in exactly the same cases.
 Theorem 1 (Compositional grounding preserves truth conditions).
@@
 Apply Lemma 1 iteratively: $\mathrm{T}(g^{i})=\mathrm{T}(g^{i+1})$ for each $i<m$, hence $\mathrm{T}(g^{0})=\mathrm{T}(g^{m})=\mathrm{T}(l^{m})$. ∎
-Proof intuition.
-Iterate Lemma 1 down the stack and every grounding step preserves the truth set, so the composed grounding does too.
-Remark 1 (Non-uniqueness (and why that is acceptable)).
-Grounding is generally non-unique because the witness completions $o_{p}$ in Lemma 1 need not be unique.
-For Temporal Gap questions, what matters is that some base-layer grounded conjunction with the same truth set exists.
-Different groundings correspond to different decompositions of the same higher-layer content into base-level ingredients.
 4 Stack-Time Semantics Module
-This section is designed as a reusable addendum to Stack Theory. It defines induced layer time, window semantics, and temporal lifting operators in a way that can be cited by future work.
+This section defines induced layer time, window semantics, and temporal lifting operators as a reusable addendum to Stack Theory.
 4.1 Objective Time and Layer Time as a Quotient
 Definition 7 (Objective time and trajectories).
-Objective time is represented by a discrete index $t\in\mathbb{N}$. An objective trajectory is a function $\tau:\mathbb{N}\to\Phi$ such that
-$\tau(t{+}1)\neq\tau(t)\quad\text{for all }t\in\mathbb{N}.$
+Objective time is represented by a discrete index $t\in\mathbb{N}$. An objective trajectory is a function $\tau:\mathbb{N}\to\Phi$ such that $\tau(t{+}1)\neq\tau(t)$ for all $t\in\mathbb{N}$.
 Intuition.
-Objective time is a step counter for objective change. If nothing changes then there is no next objective index. Two events (statements) are simultaneous when they occur at the same objective index.
+Objective time is a step counter for objective change. If nothing changes then there is no next objective index. Two events are simultaneous when they occur at the same objective index.
 Definition 8 (Encoding map).
-For a vocabulary $\mathfrak{v}\subseteq P$ and a state $\phi\in\Phi$, define
-$\mathrm{Enc}_{\mathfrak{v}}(\phi):=\{\,p\in\mathfrak{v}:\phi\in p\,\}\subseteq\mathfrak{v}.$
+For a vocabulary $\mathfrak{v}\subseteq P$ and a state $\phi\in\Phi$, define $\mathrm{Enc}_{\mathfrak{v}}(\phi):=\{\,p\in\mathfrak{v}:\phi\in p\,\}\subseteq\mathfrak{v}$.
 Intuition.
-Given a state and a vocabulary, the encoding collects all programs in that vocabulary that are true at that state.
-This gives the most detailed statement you can write in that vocabulary that still holds at that state.
+Given a state and a vocabulary, the encoding collects all programs in that vocabulary true at that state—the most detailed true statement in that vocabulary.
 Lemma 2 (Encoding yields the maximal true statement).
@@
 Moreover, if $l\in\mathcal{L}_{\mathfrak{v}}$ is true at $\phi$, then $l\subseteq\mathrm{Enc}_{\mathfrak{v}}(\phi)$.
-Proof.
-Since $\phi\in p$ for every $p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)$, I have $\phi\in\bigcap_{p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)}p=\mathrm{T}(\mathrm{Enc}_{\mathfrak{v}}(\phi))$, so the intersection is nonempty and $\mathrm{Enc}_{\mathfrak{v}}(\phi)\in\mathcal{L}_{\mathfrak{v}}$.
-If $l$ is true at $\phi$, then $\phi\in p$ for all $p\in l$, hence $p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)$ for all $p\in l$, so $l\subseteq\mathrm{Enc}_{\mathfrak{v}}(\phi)$.
-∎
-Proof intuition.
-Every program in $\mathrm{Enc}_{\mathfrak{v}}(\phi)$ contains $\phi$, so their intersection contains $\phi$ and any statement true at $\phi$ can only use such programs.
+Proof. See Supplementary Material. ∎
 Definition 9 (Induced layer trajectory and layer time).
-Given an objective trajectory $\tau$ and a vocabulary $\mathfrak{v}$, define the induced layer trajectory
-$\tau^{(\mathfrak{v})}(t):=\mathrm{Enc}_{\mathfrak{v}}(\tau(t))\in\mathcal{L}_{\mathfrak{v}}.$
+Given an objective trajectory $\tau$ and a vocabulary $\mathfrak{v}$, define the induced layer trajectory $\tau^{(\mathfrak{v})}(t):=\mathrm{Enc}_{\mathfrak{v}}(\tau(t))\in\mathcal{L}_{\mathfrak{v}}$.
 Define the layer-tick map $\kappa_{\mathfrak{v}}:\mathbb{N}\to\mathbb{N}$ by $\kappa_{\mathfrak{v}}(0)=0$ and, for $t>0$,
 $\kappa_{\mathfrak{v}}(t):=\begin{cases}\kappa_{\mathfrak{v}}(t{-}1)&\text{if }\tau^{(\mathfrak{v})}(t)=\tau^{(\mathfrak{v})}(t{-}1),\\ \kappa_{\mathfrak{v}}(t{-}1)+1&\text{otherwise.}\end{cases}$
-Define an equivalence relation on objective times by $t\approx_{\mathfrak{v}}t^{\prime}$ iff $\kappa_{\mathfrak{v}}(t)=\kappa_{\mathfrak{v}}(t^{\prime})$.
-The corresponding layer time is the quotient set $\mathbb{N}/\!\approx_{\mathfrak{v}}$.
-Each equivalence class is a maximal contiguous block of objective times on which the induced macrostatement $\tau^{(\mathfrak{v})}$ is constant.
+Define $t\approx_{\mathfrak{v}}t^{\prime}$ iff $\kappa_{\mathfrak{v}}(t)=\kappa_{\mathfrak{v}}(t^{\prime})$.
+The layer time is $\mathbb{N}/\!\approx_{\mathfrak{v}}$: each equivalence class is a maximal contiguous block of objective times on which $\tau^{(\mathfrak{v})}$ is constant.
 Intuition.
-For each objective time step, write down the encoding at that step.
-If the encoding stays the same for several steps, treat those steps as one higher level moment.
-The tick counter increases only when the encoding changes, so higher level time advances only when something visible at that level changes.
+For each objective time step, write down the encoding.
+If it stays the same for several steps, treat them as one higher level moment.
+Layer time advances only when something visible at that level changes.
 Proposition 1 (Layer time can be arbitrarily sparse).
 For any $N\in\mathbb{N}$ there exist $\Phi,\mathfrak{v},\tau$ such that the induced trajectory $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps.
-The objective-time gap between successive layer changes can be arbitrarily large.
-Proof.
-Let $\Phi=\{\phi_{t}:t\in\mathbb{N}\}$.
-Let $p:=\{\phi_{0},\phi_{1},\dots,\phi_{N}\}$ and take $\mathfrak{v}=\{p\}$.
-Define $\tau(t)=\phi_{t}$ for all $t\in\mathbb{N}$.
-Then $\tau$ is an objective trajectory because $\tau(t{+}1)\neq\tau(t)$ for all $t$.
-For all $t\leq N$ I have $\tau(t)\in p$, so $\tau^{(\mathfrak{v})}(t)=\{p\}$.
-I also have $\tau(N{+}1)\notin p$, so $\tau^{(\mathfrak{v})}(N{+}1)=\emptyset$.
-Hence $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps.
-∎
-Proof intuition.
-Choose a vocabulary that cannot distinguish many distinct objective microstates. The induced macrostate then stays constant across many objective changes, so layer time barely advances.
+Proof. See Supplementary Material. Choose a vocabulary that cannot distinguish many distinct objective microstates; the macrostate stays constant. ∎
 4.2 Time and Lifting Operators
-A subjective moment is commonly modelled as having nonzero temporal extent (the “specious present”) [undefr, undefaa].
-To speak precisely about satisfaction within such a window, I introduce windowing maps and their induced temporal operators.
-Definition 10 (Windowing maps).
-A windowing map is a function $\mathrm{W}:\mathbb{N}\to\mathcal{P}(\mathbb{N})$ such that $\mathrm{W}(t)$ is finite and nonempty for every $t$.
-For horizon $\Delta\in\mathbb{N}$ and stride $s\in\mathbb{N}_{>0}$, the sliding window map is
-$\mathrm{W}_{\Delta,s}(t)=\{st,st{+}1,\dots,st{+}\Delta\}.$
-Intuition.
-A windowing map picks, for each subjective index, a small set of objective time steps that belong to that moment.
-The parameter $\Delta$ controls how far the window extends, so the window contains $\Delta{+}1$ time steps.
-The parameter $s$ controls how far you shift the window each time.
+A subjective moment is commonly modelled as having nonzero temporal extent (the “specious present”) [19, 28]. I introduce temporal operators to formalise satisfaction within such a window.
 Definition 10 (Ingredient-wise vs co-instantiated satisfaction).
-Fix a windowing map $\mathrm{W}$, an objective trajectory $\tau$, and a subjective index $t$. For a program $p\subseteq\Phi$, define
+Fix a windowing map $\mathrm{W}$, an objective trajectory $\tau$, and a window index $t$. For a program $p\subseteq\Phi$, define
 $\mathrm{Occur}_{\mathrm{W}}(p,\tau,t)\iff\exists u\in\mathrm{W}(t)\ \big(\tau(u)\in p\big).$
@@
 Intuition.
-Ingredient-wise means every ingredient becomes true at least once somewhere in the window. Co-instantiated means there is some time step in the window where all ingredients are true together.
+Ingredient-wise means every ingredient becomes true at least once somewhere in the window. Co-instantiated means there is a time step in the window where all ingredients are true together.
 Remark 1 (Quantifier structure).
-For conjunctive $l$, ingredient-wise satisfaction is of the form $\forall p\in l\ \exists u\in\mathrm{W}(t)$, while co-instantiation is $\exists u\in\mathrm{W}(t)\ \forall p\in l$.
-The Temporal Gap is precisely the possible separation between these quantifier orders.
+For conjunctive $l$, ingredient-wise satisfaction is $\forall p\in l\ \exists u\in\mathrm{W}(t)$, while co-instantiation is $\exists u\in\mathrm{W}(t)\ \forall p\in l$.
+The Temporal Gap amounts to whether conscious unity requires ingredient-wise satisfaction or co-instantiation.
 Definition 11 (Window environments and window trajectories).
@@
 Define the window environment $\Phi^{\Delta}:=\Phi^{\{0,1,\dots,\Delta\}}\cong\Phi^{\Delta+1}$ whose elements are windows $\sigma=(\phi_{0},\dots,\phi_{\Delta})$.
-Given stride $s$ and an objective trajectory $\tau$, define the induced window trajectory
-$\tau^{\Delta,s}(t):=\big(\tau(st),\tau(st{+}1),\dots,\tau(st{+}\Delta)\big)\in\Phi^{\Delta}.$
+Given an objective trajectory $\tau$, define the induced window trajectory $\tau_{\Delta}(t):=\big(\tau(t),\tau(t{+}1),\dots,\tau(t{+}\Delta)\big)\in\Phi^{\Delta}$.
 Intuition.
-Treat a whole window of $\Delta{+}1$ consecutive states as a single new state. This lets you test lifted statements one window at a time.
+Treat a whole window of $\Delta{+}1$ consecutive states as a single new state.
 Definition 12 (Temporal lifts of programs).
@@
 Intuition.
-$\Diamond_{\Delta}p$ means $p$ is true at least once somewhere in the window. $\Box_{\Delta}p$ means $p$ is true at every time step in the window.
+$\Diamond_{\Delta}p$ means $p$ is true at least once in the window. $\Box_{\Delta}p$ means $p$ is true at every step in the window.
 Definition 13 (Lifted statements).
@@
 Intuition.
-To lift a statement, lift each ingredient and then require all lifted ingredients to hold for the window.
-With $\Diamond_{\Delta}$ this means every ingredient shows up somewhere in the window.
-With $\Box_{\Delta}$ this means every ingredient stays true throughout the window.
-Remark 3 (Internalising temporal semantics as an abstraction layer).
-The window environment $\Phi^{\Delta}$ and the lifted vocabulary $\{\Diamond_{\Delta}p:p\in\mathfrak{v}\}$ (or $\{\Box_{\Delta}p\}$) together define a new abstraction layer in the native Stack-Theory sense.
-This is the formal bridge between “over time” constraints and pointwise Stack-Theory semantics.
+To lift a statement, lift each ingredient and require all lifted ingredients to hold. With $\Diamond_{\Delta}$ this means every ingredient shows up somewhere in the window.
+Remark 2 (Temporal semantics as an abstraction layer).
+The window environment $\Phi^{\Delta}$ and the lifted vocabulary $\{\Diamond_{\Delta}p:p\in\mathfrak{v}\}$ define a new abstraction layer in the native Stack-Theory sense, bridging “over time” constraints and pointwise semantics.
+The compilation of history-dependent (pathwise) abstractors into pointwise abstractors on augmented state is detailed in the Supplementary Material.
 4.3 The Algebraic Core
-The following commutation and non-commutation facts are elementary, but spelling them out in Stack-Theory form turns the Temporal Gap into a concrete architectural question.
 Theorem 2 (Universal lift commutes with conjunction).
-For any statement $l\in\mathcal{L}_{\mathfrak{v}}$ and any $\Delta\in\mathbb{N}$,
-$\Box_{\Delta}\mathrm{T}(l)\,=\,\mathrm{T}(\Box_{\Delta}l).$
-Proof.
-Fix any window $\sigma\in\Phi^{\Delta}$. Then
-$\sigma\in\mathrm{T}(\Box_{\Delta}l)$ $\iff\forall p\in l,\ \sigma\in\Box_{\Delta}p$
-$\iff\forall p\in l,\ \forall k\leq\Delta,\ \sigma(k)\in p$
-$\iff\forall k\leq\Delta,\ \sigma(k)\in\bigcap_{p\in l}p$
-$\iff\sigma\in\Box_{\Delta}\mathrm{T}(l).$
-Proof intuition.
-$\Box_{\Delta}$ is “for all positions in the window”. Since $\forall$ distributes over conjunction, $\Box_{\Delta}$ commutes with $\wedge$.
+For any statement $l\in\mathcal{L}_{\mathfrak{v}}$ and any $\Delta\in\mathbb{N}$, $\Box_{\Delta}\mathrm{T}(l)=\mathrm{T}(\Box_{\Delta}l)$.
+Proof. See Supplementary Material. $\Box_{\Delta}$ is “for all positions”; since $\forall$ distributes over conjunction, $\Box_{\Delta}$ commutes with $\wedge$. ∎
 Theorem 3 (Existential lift does not commute with conjunction).
@@
 Remark 3 (Degenerate equality cases).
-If $\Delta=0$ (a window is a single instant) or $\lvert l\rvert\leq 1$ (no nontrivial conjunction), then
-$\Diamond_{\Delta}\mathrm{T}(l)\,=\,\mathrm{T}(\Diamond_{\Delta}l).$
-Thus strict separation requires $\Delta\geq 1$ and at least two ingredients.
+If $\Delta=0$ or $\lvert l\rvert\leq 1$, then $\Diamond_{\Delta}\mathrm{T}(l)=\mathrm{T}(\Diamond_{\Delta}l)$.
+Strict separation requires $\Delta\geq 1$ and at least two ingredients.
+A sufficient condition for commutation along a trajectory (within-window persistence) is given in the Supplementary Material.
 Remark 4 (Reading the theorem back into window semantics).
 In the sliding-window case, Definition 10 corresponds to
-$\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(l,\tau,t)$ $\iff\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}l),$
-$\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(l,\tau,t)$ $\iff\tau^{\Delta,s}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
-Thus Theorem 3 is exactly the claim that ingredient-wise satisfaction is strictly weaker than co-instantiation in general.
-4.4 When does $\Diamond$ commute with conjunction?
-Definition 15 (Within-window persistence along a trajectory).
-Fix $\Delta\in\mathbb{N}$ and an objective trajectory $\tau:\mathbb{N}\to\Phi$.
-A program $p\subseteq\Phi$ is $(\tau,\Delta)$-persistent if for all $t\in\mathbb{N}$ and all indices $k\leq k^{\prime}\leq\Delta$,
-$\tau(t+k)\in p\ \Longrightarrow\ \tau(t+k^{\prime})\in p.$
-Intuition.
-Along the trajectory, if $p$ becomes true at some point inside a $\Delta$ window, then $p$ stays true for the rest of that same window.
-Proposition 2 (A sufficient condition for co-instantiation along a trajectory).
-Fix $\Delta\in\mathbb{N}$ and an objective trajectory $\tau$.
-Let $l\in\mathcal{L}_{\mathfrak{v}}$ be a statement such that every ingredient program $p\in l$ is $(\tau,\Delta)$-persistent.
-Then for all $t$,
-$\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l)\quad\Longrightarrow\quad\tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
-Proof.
-Assume $\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l)$.
-Then for each $p\in l$ there exists $k_{p}\leq\Delta$ with $\tau(t+k_{p})\in p$.
-Let $k^{\ast}:=\max_{p\in l}k_{p}$.
-By $(\tau,\Delta)$-persistence of each $p$, from $\tau(t+k_{p})\in p$ and $k_{p}\leq k^{\ast}\leq\Delta$ I obtain $\tau(t+k^{\ast})\in p$ for all $p\in l$.
-Hence $\tau(t+k^{\ast})\in\bigcap_{p\in l}p=\mathrm{T}(l)$, so $\tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(l)$.
-∎
-Proof intuition.
-With persistence, once each ingredient appears it remains true. Choosing the latest appearance time yields a single instant where all ingredients overlap.
-4.5 Finite-horizon temporal semantics
-Finite-horizon (history-dependent) temporal semantics can itself be treated as an abstraction layer.
-A horizon-$\Delta$ pathwise abstractor can be compiled into a pointwise abstractor on an augmented state, so Stack-Theory tools apply without modifying the underlying semantics.
-Definition 16 (Pointwise and horizon-$\Delta$ pathwise abstractors).
-Let $\Psi$ be any higher-level state space. A pointwise abstractor is a function $A:\Phi\to\Psi$. A pathwise abstractor of horizon $\Delta$ is a function $A^{(\Delta)}:\Phi^{\Delta}\to\Psi$.
-Intuition.
-A pointwise abstractor looks only at the current state. A horizon $\Delta$ pathwise abstractor looks at the recent $\Delta{+}1$ states, so it can use short term history.
-Theorem 4 (History-state (state augmentation) compilation).
-Fix $\Delta\in\mathbb{N}$ and let $A^{(\Delta)}:\Phi^{\Delta}\to\Psi$ be a pathwise abstractor.
-Define the augmented environment $\widetilde{\Phi}:=\Phi^{\Delta}$ and the pointwise abstractor $\widetilde{A}:\widetilde{\Phi}\to\Psi$ by $\widetilde{A}(\sigma):=A^{(\Delta)}(\sigma)$.
-Then for every objective trajectory $\tau:\mathbb{N}\to\Phi$ and every $t$, with augmented trajectory $\widetilde{\tau}(t):=\tau^{\Delta,1}(t)\in\widetilde{\Phi}$ I have
-$\widetilde{A}(\widetilde{\tau}(t))=A^{(\Delta)}(\tau^{\Delta,1}(t)).$
-Proof.
-Immediate from the definitions of $\widetilde{A}$ and $\widetilde{\tau}$. ∎
-Proof intuition.
-Encode the last $\Delta$ microstates into an augmented state so any horizon-$\Delta$ (history-dependent) abstractor becomes pointwise on the augmented state.
-5 StrongSync and WeakSync Postulates
+$\mathrm{Occur}_{\mathrm{W}_{\Delta}}(l,\tau,t)$ $\iff\tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}l),$
+$\mathrm{CoInst}_{\mathrm{W}_{\Delta}}(l,\tau,t)$ $\iff\tau_{\Delta}(t)\in\Diamond_{\Delta}\mathrm{T}(l).$
+Theorem 3 is exactly the claim that ingredient-wise satisfaction is strictly weaker than co-instantiation in general.
+5 Chord and Arpeggio Postulates
 The above algebra does not decide which window semantics corresponds to phenomenality.
-It clarifies what must be decided.
-I’ll now state StrongSync and WeakSync in terms of a higher-layer moment statement and its base-layer grounding across a stack state.
+I now state Chord and Arpeggio in terms of occurrence and co-instantiation, with a higher-layer moment statement and its base-layer grounding.
 Assumption 1 (Subjective moments live at some layer).
@@
 Definition 14 (Phenomenal realisation predicate).
-Let $\mathrm{PhenReal}(l^{m},\tau,t)$ mean “moment statement $l^{m}$ is phenomenally realised at (higher-layer) time $t$ along objective trajectory $\tau$.” I treat $\mathrm{PhenReal}$ as primitive, constrained by StrongSync/WeakSync.
+Let $\mathrm{PhenReal}(l^{m},\tau,t)$ mean “moment statement $l^{m}$ is phenomenally realised at (higher-layer) time $t$ along objective trajectory $\tau$.” I treat $\mathrm{PhenReal}$ as primitive, constrained by Chord/Arpeggio.
 Intuition.
-$\mathrm{PhenReal}$ names the claim that moment statement $l^{m}$ is experienced at subjective time $t$ along the run $\tau$.
-Here $l^{m}$ can include first, second, and third order selves and causal identities for objects and properties [undeff].
+$\mathrm{PhenReal}$ names the claim that moment $l^{m}$ is experienced at subjective time $t$ along run $\tau$.
+Here $l^{m}$ can include first, second, and third order selves and causal identities for objects and properties [7].
 Definition 15 (Base grounding of a moment statement).
@@
 I call $g^{0}$ the grounded conjunction of the moment statement.
-Intuition.
-$g^{0}$ is the low level conjunction that corresponds to the higher level moment statement $l^{m}$.
-It is built by expanding $l^{m}$ down to the base layer.
-It is true in exactly the same situations as $l^{m}$.
-Definition 19 (StrongSync (objective co-instantiation required)).
-Fix a windowing map for the phenomenal layer, typically the sliding window $\mathrm{W}_{\Delta,s}$ of horizon $\Delta$ and stride $s$ (Section 4.2). StrongSync postulates that for all $l^{m},\tau,t$,
-$\mathrm{PhenReal}(l^{m},\tau,t)$ $\Rightarrow\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)$
-$\text{(equivalently }\tau^{\Delta,s}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\text{).}$
+Definition 16 (Chord (objective co-instantiation required)).
+Fix a windowing map for the phenomenal layer, typically the sliding window $\mathrm{W}_{\Delta}$ of horizon $\Delta$ (Section 4.2). Chord postulates that for all $l^{m},\tau,t$,
+$\mathrm{PhenReal}(l^{m},\tau,t)$ $\Rightarrow\mathrm{CoInst}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t)$
+$\text{(equivalently }\tau_{\Delta}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\text{).}$
 If the moment is experienced, then within the corresponding subjective window there exists an objective time-slice at which the entire grounded conjunction holds.
 Intuition.
-StrongSync says that if the moment is experienced, then within the window there is at least one objective time step where all grounded ingredients are true together.
-Definition 20 (WeakSync (objective smearing permitted)).
-Fix the same windowing map $\mathrm{W}_{\Delta,s}$ as in Definition 19. WeakSync postulates the following.
+Chord says: if the moment is experienced, then within the window there is at least one objective time step where all grounded ingredients are true together.
+Definition 17 (Arpeggio (objective smearing permitted)).
+Fix the same windowing map $\mathrm{W}_{\Delta}$ as in Definition 16. Arpeggio postulates the following.
   1. 1.
     Ingredient-wise necessity. For all $l^{m},\tau,t$,
-    $\mathrm{PhenReal}(l^{m},\tau,t)\Rightarrow\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)$
-    $\text{(equivalently}\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\text{).}$
+    $\mathrm{PhenReal}(l^{m},\tau,t)\Rightarrow\mathrm{Occur}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t)$
+    $\text{(equivalently}\tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\text{).}$
   2. 2.
     Smearing permitted. There exist $l^{m},\tau,t$ such that
-    $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\mathrm{Occur}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t)\wedge\neg\mathrm{CoInst}_{\mathrm{W}_{\Delta,s}}(g^{0},\tau,t),$
-    equivalently $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\tau^{\Delta,s}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\wedge\tau^{\Delta,s}(t)\notin\Diamond_{\Delta}\mathrm{T}(g^{0})$.
-In plain English this allows a moment to be experienced without any instant realising the full grounded conjunction.
+    $\mathrm{PhenReal}(l^{m},\tau,t)\wedge\mathrm{Occur}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t)\wedge\neg\mathrm{CoInst}_{\mathrm{W}_{\Delta}}(g^{0},\tau,t).$
+This allows a moment to be experienced without any instant realising the full grounded conjunction.
 Intuition.
-WeakSync says that if the moment is experienced, then each grounded ingredient must occur somewhere in the window.
-It allows that there may be no single objective time step where they are all true together.
+Arpeggio says: if the moment is experienced, each grounded ingredient must occur somewhere in the window, but there need not be a single time step where they are all true together.
 6 Architectural Consequences
-A polycomputer simultaneously executes multiple functions using the same parts or physical substrate [undefk].
-A liquid brain such as an ant colony does not rely on a persistent structure to compute, but instead computes through movement [undefx].
-In contrast, a “solid brain” is one that has a persistent structure that can support, for example, a bio-electric abstraction layer with synchronous broadcast [undeff] (like a network of telegraph cables).
-I now translate “single-thread sequential emulation” versus “synchronous, solid-brained, polycomputational multi-contributor realisation” into a minimal architectural invariant with a bound on simultaneous causal contribution.
-The Temporal Gap determines whether that bound is consciousness-relevant.
+A polycomputer simultaneously executes multiple functions using the same parts or physical substrate [13].
+A liquid brain such as an ant colony computes through movement rather than persistent structure [25].
+In contrast, a “solid brain” has persistent structure supporting, e.g., a bio-electric abstraction layer with synchronous broadcast [7].
+I now translate “single-thread sequential emulation” versus “synchronous polycomputational realisation” into a minimal architectural invariant bounding simultaneous causal contribution.
+The Temporal Gap asks whether that bound is consciousness-relevant (and whether conscious unity requires causal exchange between co-instantiated ingredients, explored in [10]).
 6.1 Concurrency Capacity as an Invariant
 Definition 18 (Contributor model).
-Fix $n\geq 2$ potential contributors and write $n]:=\{1,\dots,n\}$. Define a base environment of microstates
-$\Phi:=X\times\mathcal{P}([n]),$
-where $x\in X$ is a “data” component and $A\subseteq[n]$ is the set of currently active contributors at that objective instant. For each $i\in[n]$ define the program
-$p_{i}:=\{(x,A)\in\Phi:i\in A\}.$
+Fix $n\geq 2$ potential contributors and write $n]:=\{1,\dots,n\}$.
+Define a base environment $\Phi:=X\times\mathcal{P}([n])$, where $x\in X$ is a “data” component and $A\subseteq[n]$ is the set of currently active contributors.
+For each $i\in[n]$ define $p_{i}:=\{(x,A)\in\Phi:i\in A\}$.
 Let $l_{n}:=\{p_{1},\dots,p_{n}\}$ and note $\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$.
 Intuition.
-The contributor model records which contributors are active at each time step.
-A conjunction that mentions many contributors can only be true when all of those contributors are active at the same time step.
+The contributor model records which contributors are active at each time step. A conjunction mentioning many contributors can only be true when all are active simultaneously.
 Definition 19 (Concurrency capacity).
-An architecture has concurrency capacity $c\in\{1,\dots,n\}$ if its admissible trajectories $\tau:\mathbb{N}\to\Phi$ satisfy (writing $\tau(t)=(x_{t},A_{t})$ for the data and contributor-set components)
-$\lvert A_{t}\rvert\leq c\quad\text{for all objective times }t.$
-The case $c=1$ formalises a sequential single-thread update model (one active contributor per tick).
-The case $c\geq n$ formalises synchronous polycomputing at the relevant scale (all contributors may be active together).
+An architecture has concurrency capacity $c\in\{1,\dots,n\}$ if its admissible trajectories satisfy $\lvert A_{t}\rvert\leq c$ for all objective times $t$ (writing $\tau(t)=(x_{t},A_{t})$).
+The case $c=1$ formalises a sequential single-thread model; $c\geq n$ formalises synchronous polycomputing at the relevant scale.
 Intuition.
-Capacity $c$ is the maximum number of contributors that can be active at one time step.
-If the grounded content needs more than $c$ contributors active together, then StrongSync cannot hold on that architecture.
+Capacity $c$ is the maximum number of contributors active at one time step. If the grounded content needs more than $c$ active together, co-instantiation cannot hold.
 Theorem 4 (Synchronous threshold for co-instantiation).
-Fix $n\geq 2$ and content statement $l_{n}$. If an architecture has concurrency capacity $c<n$, then along every admissible trajectory $\tau$ and for every horizon $\Delta$ I have
-$\tau^{\Delta,1}(t)\notin\Diamond_{\Delta}\mathrm{T}(l_{n})\quad\text{for all }t.$
-However, if the architecture is such that every contributor is activated at least once within some horizon-$\Delta$ window (i.e.
-for some $t$ and all $i$ there exists $k_{i}\leq\Delta$ with $i\in A_{t+k_{i}}$), then
-$\tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}l_{n}).$
+Fix $n\geq 2$ and content statement $l_{n}$.
+If an architecture has concurrency capacity $c<n$, then along every admissible trajectory $\tau$ and for every horizon $\Delta$, $\tau_{\Delta}(t)\notin\Diamond_{\Delta}\mathrm{T}(l_{n})$ for all $t$.
+However, if every contributor is activated at least once within some horizon-$\Delta$ window, then $\tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}l_{n})$.
 Thus the architecture admits ingredient-wise satisfaction but forbids objective co-instantiation.
 Proof.
-If $c<n$, then at any time $t$ I have $\lvert A_{t}\rvert\leq c<n$, so $A_{t}\neq[n]$.
-Hence $\tau(t)\notin\mathrm{T}(l_{n})=\{(x,[n]):x\in X\}$ for all $t$.
-Therefore no window contains an element of $\mathrm{T}(l_{n})$, so $\Diamond_{\Delta}\mathrm{T}(l_{n})$ is never satisfied.
-For the second claim, if each $i$ appears in $A_{t+k_{i}}$ for some $k_{i}\leq\Delta$, then the window $\sigma=\tau^{\Delta,1}(t)$ satisfies $\sigma\in\Diamond_{\Delta}p_{i}$ for all $i$, hence $\sigma\in\bigcap_{i}\Diamond_{\Delta}p_{i}=\mathrm{T}(\Diamond_{\Delta}l_{n})$.
+If $c<n$, then $\lvert A_{t}\rvert\leq c<n$ at every $t$, so $A_{t}\neq[n]$ and $\tau(t)\notin\mathrm{T}(l_{n})$ for all $t$.
+No window contains an element of $\mathrm{T}(l_{n})$, so $\Diamond_{\Delta}\mathrm{T}(l_{n})$ is never satisfied.
+For the second claim, if each $i$ appears in $A_{t+k_{i}}$ for some $k_{i}\leq\Delta$, then $\sigma=\tau_{\Delta}(t)$ satisfies $\sigma\in\Diamond_{\Delta}p_{i}$ for all $i$, hence $\sigma\in\mathrm{T}(\Diamond_{\Delta}l_{n})$.
 ∎
@@
 If at most $c$ contributors can be active, an $n$-way conjunction with $n>c$ is never true at an instant, though each ingredient can appear somewhere in a window.
-Corollary 1 (Single-thread CPU vs synchronous polycomputer (formal)).
+Corollary 1 (Single-thread CPU vs synchronous polycomputer).
 Let a moment statement $l^{m}$ have base grounding $g^{0}$ whose ingredients include a conjunctive substatement equivalent to $l_{n}$ for some $n\geq 2$.
-Fix the phenomenal windowing map $\mathrm{W}_{\Delta,s}$ used in StrongSync/WeakSync.
+Fix the phenomenal windowing map $\mathrm{W}_{\Delta}$ used in Chord/Arpeggio.
   1. 1.
-    Under StrongSync, any sequential-update architecture with capacity $c<n$ cannot phenomenally realise $l^{m}$, because $\Diamond_{\Delta}\mathrm{T}(g^{0})$ fails for the relevant grounded substatement (Theorem 5).
+    Under Chord, any sequential-update architecture with capacity $c<n$ cannot phenomenally realise $l^{m}$ (Theorem 4).
   2. 2.
-    Under WeakSync, the same sequential architecture remains a candidate in principle, because $\mathrm{T}(\Diamond_{\Delta}g^{0})$ may hold without $\Diamond_{\Delta}\mathrm{T}(g^{0})$.
+    Under Arpeggio, the same sequential architecture remains a candidate, because $\mathrm{T}(\Diamond_{\Delta}g^{0})$ may hold without $\Diamond_{\Delta}\mathrm{T}(g^{0})$.
   3. 3.
-    Any synchronous architecture with capacity $c\geq n$ can satisfy $\Diamond_{\Delta}\mathrm{T}(g^{0})$ whenever it activates the required contributors simultaneously, hence can host the moment under both StrongSync and WeakSync.
+    Any synchronous architecture with capacity $c\geq n$ can satisfy $\Diamond_{\Delta}\mathrm{T}(g^{0})$ by activating the required contributors simultaneously, hence can host the moment under both postulates.
 Remark 5 (Connection to valence-first Stack Theory).
-My valence-first application of Stack Theory treats certain phenomenal configurations as synchronised, causally efficacious conjunctions over many ingredients [undeff].
-The capacity model above isolates one minimal implementation parameter, simultaneous contribution, that controls whether such conjunctions can be instantiated at any objective instant versus only “smeared” across a window.
-7 Evidence for StrongSync
-StrongSync and WeakSync predict different empirical signatures within a short subjective window.
-Either consciousness requires a window-local episode of coordinated co-engagement (objective co-instantiation), or it can tolerate mere ingredient-wise activation distributed across the window.
+My valence-first application of Stack Theory treats certain phenomenal configurations as synchronised, causally efficacious conjunctions over many ingredients [7].
+The capacity model isolates one minimal implementation parameter—simultaneous contribution—controlling whether such conjunctions can be instantiated at any objective instant versus only “smeared” across a window.
+7 Evidence for Chord
+Chord and Arpeggio predict different empirical signatures within a short subjective window.
+Either consciousness requires window-local coordinated co-engagement, or it can tolerate mere ingredient-wise activation distributed across the window.
 Several classic paradigms are more naturally described by the coordination view.
-In masking, conscious perception has been reported to correlate with transient long-range gamma phase synchrony across cortical areas, even when local activity changes are similar across seen versus unseen stimuli [undeft].
-In non-REM sleep, TMS-EEG responses can become strong yet local, failing to propagate, consistent with a breakdown of effective connectivity when consciousness fades [undefs].
-More generally, perturbational measures such as PCI track level of consciousness across wake/sleep/anesthesia and disorders of consciousness, consistent with a requirement for integrated, differentiated system-level dynamics [undefm].
-Taken together, these findings do not prove StrongSync, but they do make the existential permissiveness of WeakSync much less compelling.
-Conscious level appears to covary with the presence (and disappearance with the collapse) of a temporally coordinated, system-level regime.
-Hence when I draw architectural conclusions below, I treat StrongSync as the default working hypothesis.
-This step remains defeasible.
+In masking, conscious perception correlates with transient long-range gamma phase synchrony across cortical areas, even when local activity is similar across seen versus unseen stimuli [21].
+In non-REM sleep, TMS-EEG responses become strong yet local, failing to propagate, consistent with breakdown of effective connectivity when consciousness fades [20].
+Perturbational measures such as PCI track level of consciousness across wake/sleep/anesthesia and disorders of consciousness, consistent with a requirement for integrated, differentiated dynamics [15].
+These findings do not prove Chord, but they make the existential permissiveness of Arpeggio less compelling.
+Conscious level covaries with the presence of a temporally coordinated, system-level regime.
+I therefore treat Chord as the default working hypothesis; this step remains defeasible.
 8 Attribution and Implications
-The machine consciousness context puts special weight on attribution/measurement under uncertainty and implementability constraints.
-The Temporal Gap formalism makes a concrete measurement target of the gap between ingredient coverage and co-instantiation for a grounded conjunction.
-8.1 Instrumentable separation metrics
-Fix a grounded conjunction $g^{0}$ and an objective trajectory $\tau$ (possibly generated under controlled conditions).
-Define the minimal horizon needed for ingredient coverage versus co-instantiation:
-$w_{\mathrm{ing}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau^{\Delta,1}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\},$
-$w_{\mathrm{co}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau^{\Delta,1}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\}.$
-I use the convention $\inf\emptyset=\infty$ (so these live in $\mathbb{N}\cup\{\infty\}$). The strict Temporal Gap pattern is $w_{\mathrm{ing}}<w_{\mathrm{co}}$ (or $w_{\mathrm{co}}=\infty$).
-These are, in principle, instrumentable when the base-layer ingredients are defined in terms of architectural state variables (e.g.
-active contributor sets).
-In the broader context of AI consciousness assessment, this kind of architectural metric aligns with the view that behaviour alone may be insufficient and that concrete indicator properties should be checked against candidate theories [undefl].
-8.2 Contrast-class experiment
-Construct paired systems that are behaviourally matched at a reporting layer (same task performance and report interface) but differ in base-layer concurrency capacity:
-  1. 1.
-    Synchronous substrate: implement a bounded-latency global tick and allow high concurrency at the relevant scale (capacity $c\geq n$).
-  2. 2.
-    Sequential substrate: implement the same macro update rule but enforce small $c$ by serialising contributor updates.
-Then select candidate contents $l^{m}$ whose groundings include large conjunctions (Section 6), and measure $w_{\mathrm{ing}}$ and $w_{\mathrm{co}}$.
-StrongSync predicts that phenomenality (if present) tracks bounded $w_{\mathrm{co}}$ at the relevant grounding layer.
-WeakSync predicts it may track $w_{\mathrm{ing}}$ alone.
-8.3 Normative significance under uncertainty
-If StrongSync is even plausible, then behavioural equivalence between a synchronous system and a sequential emulation does not settle whether they are phenomenally equivalent.
+The Temporal Gap is the question of whether consciousness is affected by the gap between ingredient coverage and co-instantiation for a grounded conjunction.
+Fix a grounded conjunction $g^{0}$ and an objective trajectory $\tau$. Define the minimal horizon needed for ingredient coverage versus co-instantiation:
+$w_{\mathrm{ing}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau_{\Delta}(t)\in\mathrm{T}(\Diamond_{\Delta}g^{0})\},$
+$w_{\mathrm{co}}(g^{0},\tau)$ $:=\inf\{\Delta\in\mathbb{N}:\exists t\ \tau_{\Delta}(t)\in\Diamond_{\Delta}\mathrm{T}(g^{0})\}.$
+The strict Temporal Gap pattern is $w_{\mathrm{ing}}<w_{\mathrm{co}}$ (or $w_{\mathrm{co}}=\infty$).
+These are instrumentable when base-layer ingredients are defined in terms of architectural state variables.
+This kind of architectural metric aligns with the view that behaviour alone may be insufficient for consciousness assessment and that concrete indicator properties should be checked against candidate theories [14].
+A concrete test follows: construct paired systems that are behaviourally matched at a reporting layer but differ in base-layer concurrency capacity (one synchronous with $c\geq n$, one sequential with small $c$).
+Select candidate contents whose groundings include large conjunctions, and measure $w_{\mathrm{ing}}$ and $w_{\mathrm{co}}$.
+Chord predicts that phenomenality tracks bounded $w_{\mathrm{co}}$; Arpeggio predicts it may track $w_{\mathrm{ing}}$ alone.
+If Chord is even plausible, behavioural equivalence between a synchronous system and a sequential emulation does not settle phenomenal equivalence.
 This creates an ethically salient zone of reasonable disagreement about moral status, particularly when a system could have valenced experience.
-A precautionary stance toward sentience under uncertainty has been defended in nearby contexts [undefi].
-Within my framework, one concrete precautionary recommendation is to treat large Temporal-Gap regimes ($w_{\mathrm{co}}\gg w_{\mathrm{ing}}$ or $w_{\mathrm{co}}=\infty$) as a high-uncertainty region for attribution, motivating conservative deployment and experimentation policies until the StrongSync/WeakSync question is better resolved.
+A precautionary stance toward sentience under uncertainty has been defended in nearby contexts [11].
+One concrete recommendation is to treat large Temporal-Gap regimes ($w_{\mathrm{co}}\gg w_{\mathrm{ing}}$ or $w_{\mathrm{co}}=\infty$) as a high-uncertainty region for attribution, motivating conservative deployment policies until the Temporal Gap is better resolved.
 9 Conclusion
-I provided a Stack-Theory-compatible Stack-Time Semantics addendum and used it to sharpen the Temporal Gap into a precise algebraic and architectural question.
-Compositional grounding (Theorem 1) ensures that any higher-layer phenomenal statement has a base-layer grounded conjunction with identical truth conditions.
-Induced layer time partitions objective time into maximal constant macro-blocks (a change-point quotient) and can be arbitrarily sparse (Proposition 1), so subjective moments need not correspond to single objective instants.
-Against that backdrop, the Temporal Gap becomes the non-commutation pattern
+I provided a Stack-Theory-compatible Stack-Time Semantics addendum and used it to make the Temporal Gap into a measurable algebraic and architectural question.
+Compositional grounding (Theorem 1) ensures any higher-layer phenomenal statement has a base-layer grounded conjunction with identical truth conditions.
+Induced layer time partitions objective time into maximal constant macro-blocks and can be arbitrarily sparse (Proposition 1), so subjective moments need not correspond to single objective instants.
+The Temporal Gap becomes a question of the non-commutation pattern
 $\Diamond_{\Delta}\mathrm{T}(l)\subseteq\mathrm{T}(\Diamond_{\Delta}l),$
-with strict inclusion in general for $\Delta\geq 1$ and multi-ingredient statements (Theorem 3, Remark 4).
+with strict inclusion in general for $\Delta\geq 1$ and multi-ingredient statements (Theorem 3).
 This yields a clean separation between ingredient-wise window occurrence and objective co-instantiation.
-StrongSync and WeakSync are then competing postulates about which notion constrains phenomenality.
-Section 7 notes that conscious level often covaries with temporally coordinated, system-level dynamics (synchrony/connectivity/complexity), which tentatively favours StrongSync over WeakSync.
-Finally, I proved that a simple architectural invariant (concurrency capacity) can force the strict Temporal Gap pattern for large conjunctions (Theorem 5).
-If StrongSync is correct and if a phenomenal content’s base grounding requires a conjunction larger than a substrate’s simultaneous causal contribution can support, then that content cannot be phenomenally realised on that substrate.
-Under StrongSync, software consciousness on a strictly sequential (capacity-$1$) substrate is impossible for any content whose grounding requires two or more simultaneous contributors.
-Real machines are not perfectly sequential at the physical level.
-Thus the practical question becomes which physical concurrency scale is relevant for grounding.
-Consciousness according to Stack Theory requires a “tapestry of valence”, supported by a solid-brained polycomputing architecture, in which each co-instantiated ingredient simultaneously acts upon the others as they act on it, all linked and synchronised.
-An extreme concurrency capacity may be needed, and still concurrency is only one (important) aspect of this.
-The central conclusion remains that consciousness attribution is hardware-sensitive, and surface-level observable behavioural equivalence alone cannot wash out the Temporal Gap.
+Chord and Arpeggio are competing postulates about which notion constrains phenomenality.
+Section 7 notes that conscious level covaries with temporally coordinated dynamics, tentatively favouring Chord.
+A simple architectural invariant (concurrency capacity) can force the strict Temporal Gap pattern for large conjunctions (Theorem 4).
+Under Chord, software consciousness on a strictly sequential substrate is impossible for any content whose grounding requires two or more simultaneous contributors.
+Real machines are not perfectly sequential at the physical level, so the practical question becomes which physical concurrency scale is relevant for grounding.
+Consciousness according to Stack Theory requires a “tapestry of valence” supported by a solid-brained polycomputing architecture in which each co-instantiated ingredient simultaneously acts upon the others, all linked and synchronised (a condition formally included in Chord in [10]).
+Extreme concurrency capacity may be needed, and still concurrency is only one aspect of this.
+The central conclusion remains that consciousness attribution is hardware-sensitive, and surface-level behavioural equivalence alone cannot wash out the Temporal Gap.